### v5

# 1] organization=School of Computer Sciences , addressline=Universiti Sains Malaysia, city=11800 USM Penang, country=Malaysia

Abstract
Despite the remarkable results of deep learning in breast cancer histopathology image classification, challenges such as data imbalance and interpretability still exist and require cross-domain knowledge and collaboration among medical experts. This study proposes a breast cancer classification method using a dual-activated lightweight attention ResNet50 model, effectively addressing data imbalance and interpretability challenges. The model fuses a pre-trained deep ResNet50 and a lightweight attention mechanism to accomplish classification by embedding an attention module in layer 4 of ResNet50 and adding two fully connected layers. The fully connected network design employs LeakyReLU and ReLU activation functions.
The model outperforms SEResNet50, DensNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50 models regarding precision, accuracy, recall, F1 score, and GMean, especially in the application performance on the BreakHis dataset. In particular, the model demonstrates significant robustness and broad applicability when dealing with the unbalanced breast cancer dataset. The model has been evaluated on histopathology images at magnification factors of 40X, 100X, 200X, and 400X, achieving accuracies of 98.5%, 98.7%, 97.9%, and 94.3%, respectively. The study comprehensively assessed the model’s performance. In the later stages of training, the validated losses and accuracies change minimally, showing that the model avoids overfitting and exhibits good generalization ability. This model exhibited the fastest convergence in all laboratory experiments, even though its parameters are not the smallest. This highlights the model’s efficacy as a lightweight attention framework, showcasing its efficiency in achieving rapid convergence without compromising performance.
keywords:
[…]
1 Introduction
Breast cancer, particularly among women, has emerged as one of the leading causes of fatal diseases, with its incidence on a continuous rise, eliciting worldwide health concerns [1, 2]. Early and accurate diagnosis and treatment are crucial in reducing the mortality rate of breast cancer. However, medical practices grapple with several obstacles, including inadequate healthcare resources and human error in interpretation, which hampers accurate and prompt detection and diagnosis of breast cancer. In this context, intelligent assisted diagnosis technology has become an innovative solution in breast cancer. Recent advancements in deep learning have had a transformative impact on breast cancer diagnosis. Deep learning theory has not only surpassed the performance limitations of traditional classification and recognition algorithms but has also integrated the entire feature extraction, feature selection, and classification process into a single step. Deep learning has simplified the training process by automatically extracting and classifying features through neural networks, eliminating redundancy and improving efficiency. These developments have revolutionized breast cancer diagnosis by improving accuracy and streamlining workflows [3]. This profound learning advancement has revolutionized breast cancer diagnosis [4, 5, 6, 7]. One key innovation in this area is developing and applying Convolutional Neural Networks (CNNs), a specific class of deep learning models designed to process data with a grid-like structure, such as medical images. The concept of CNN was first introduced back in 1998 [8]. The concept describes a method for efficiently classifying image data. CNNs gradually extract abstract features of an image through multi-level convolutional and pooling operations, allowing the network to extract necessary features from different categories of images so that the different categories can be classified accurately based on those features. CNNs have three main layers: convolutional, pooling, and fully connected. These layers work together to make CNNs a powerful tool for classifying medical images. The convolutional layer captures the image’s local features, the pooling layer reduces data dimensions while preserving significant information, and the fully connected layer maps these extracted features to various output classes, As shown in Fig. 1. However, despite the significant achievements of deep learning in medical image classification, conventional CNN architectures struggle with unbalanced medical image data. To tackle this issue, In this paper, a lightweight attention mechanism based on ResNet50 [9] is introduced to address the challenge of unbalanced medical image data in conventional CNN architectures. This novel approach effectively handles unstable category distributions and accurately detects malignant breast cancer in medical images. The critical contributions of the work can be summarized as follows:
(a) Proposing a lightweight attention mechanism network that adaptively adjusts channel feature extraction to improve feature representation.
(b) Integrating the attention mechanism module and ResNet50 to form a feature-rich model, the model achieves more efficient feature extraction based on fusing the two techniques. Specifically, in the fully connected network part of the attention mechanism module, we employed LeakyReLU [10] and ReLU [11] activation functions. Leaky-ReLU was chosen to mitigate the gradient vanishing problem, while ReLU was used to enhance the image feature representation, thus better capturing multi-scale spatial patterns.
(c) In this paper, test images with different magnifications and verify that the network is invariant to image scales and performs better than other state-of-the-art methods.
(d) The model was thoroughly evaluated by analyzing loss. The results show stable performance in the model’s later training phases, minimizing overfitting and exhibiting strong generalization for unfamiliar data.
This paper is organized as follows:
Section 2 provides an overview of the current state of the art in applying deep learning to breast cancer diagnostic techniques.
Section 3 details the proposed breast cancer classification method based on dual activation lightweight attention ResNet50.
Section 4 presents the dataset and implementation details of the experimental task and compares the classification results with other state-of-the-art networks.
Section 5 comprehensively evaluates the model’s performance through loss and accuracy analysis.
Finally, Section 6 provides a summary and conclusions.
Fig. 1: Overview of Convolutional Neural Network
2 Related works
Currently, two main methods are primarily used for cancer tissue histopathology image classification: traditional manual feature extraction and deep learning-based methods. Traditional methods rely on manually extracting image features and then using classifiers such as SVM or Random Forest for recognition; however, this method has shortcomings, such as a time-consuming feature extraction process, the requirement of domain knowledge, and the challenge of securing high-quality features [12]. In contrast, methods based on deep learning extract features through models like convolutional neural networks, which is more efficient [13]. Deep learning-based cancer tissue histopathology image classification primarily encompasses three methods: CNN-based breast cancer classification, transfer learning-based breast cancer classification, and attention mechanism-based breast cancer classification. The following section delves into these techniques in detail.
2.1 CNN
CNN [8], one of the most prevalent deep learning techniques, is apt for tasks involving image processing [14]. Through convolutional, pooling, and fully connected layers, CNNs are highly effective in learning image features and classifying them [15]. In breast cancer histopathology image classification, CNN models with various structures, such as AlexNet, Inception, VGGNet, DenseNet, ResNet, etc., can be trained on large datasets to learn features and classify breast cancer.
AlexNet, a deep convolutional neural network introduced by Alex et al. in 2012 [16], marked a pivotal breakthrough in the ImageNet image classification competition, setting a landmark in deep learning and CNNs. Omonigho et al. [17] employed a modified AlexNet with augmentation techniques on the MIAS mammography database for breast cancer detection, achieving an impressive 95.70% accuracy. This surpassed traditional methods, though parameter optimization and validation on unbalanced datasets require further exploration.
The Inception model is a neural network architecture introduced by Google, which aims to solve the computation problem and several parameters in neural networks and is suitable for image classification and related tasks. The hybrid model based on CNN-Inception-V4 was used by Nazir et al. [18]. It has excellent accuracy and sensitivity, but the model’s specificity is a little inferior, and it does not consider the processing computational efficiency and resource consumption issues.
VGGNet, a deep neural network architecture developed by researchers at Oxford University, performs well in image classification tasks. However, it is unsuitable for embedded devices or resource-constrained environments because of its high computational and memory requirements due to its considerable depth and number of parameters. Nedjar et al .[19] proposed a visualization method for explaining and understanding the decision-making process of VGG19 in classifying breast biopsy images, which helps to increase the trust and acceptance of medical professionals while providing a method for evaluating and Enhancing the efficiency of CNNs.
DenseNet is a deep neural network architecture developed by Gao Huang et al. [20]. It is known for its unique connectivity pattern. Each layer is densely connected to all previous layers, thus enabling the effectiveness of information flow and gradient propagation in deep networks. Jiménez Gaona et al. [21] introduced a novel approach to breast cancer detection and classification using the DenseNet architecture, which has advantages over other pre-trained CNN methods in solving problems such as gradient vanishing and reducing the number of parameters.
ResNet, a deep convolutional neural network architecture introduced by He et al. in 2015 [9], is particularly notable for its solution to the vanishing gradient issue in deep neural networks. This enables the development of intense neural networks, leading to substantial improvements in tasks like image categorization. Khikani et al. [22] proposed an enhanced capsule network using Res2Net blocks and additional convolutional layers to extract multi-scale features, and the results showed that their model outperformed previous deep-learning methods for breast cancer classification. However, ResNet-50 is prone to problems in handling unbalanced datasets. Given the significant superiority of the ResNet50 model over other methods in terms of accuracy, this paper has decided to modify the ResNet50 model to address these specific challenges.
2.2 Transfer learning
Transfer learning applies pre-trained models from one image domain to another, effectively bridging the gap between different image domains. In the context of breast cancer image classification, this pre-trained model approach involves fine-tuning CNN models, which were initially trained on a comprehensive large-scale image data set, and then fine-tuning it to adapt to the data set specifically used for breast cancer detection. Through this step, the model is optimized to better meet the specific needs of breast cancer classification. By using the inherent ability of the pre-trained model in generalization and feature extraction, this strategy significantly improves the efficiency of training and the accuracy of classification performance [23, 24, 25, 26, 27, 28, 29].
Ismail et al. [30] used transfer learning to expedite model training. Using models like VGG16 and ResNet50, they classified breast cancer images. Their experiments confirmed that transfer learning quickens model convergence and outperforms training from scratch.
Muhammad et al. [31] also utilized pre-trained deep learning models for breast cancer detection, such as ResNet50, ResNet101, VGG16, and VGG19. After training and testing these models on a dataset of 2453 breast histopathology images, they found that ResNet50 had a superior performance with an accuracy rate of 90.2%, an Area Under the Curve (AUC) of 90.0%, and a recall rate of 94.7%.
In conclusion, considering the exceptional robustness of ResNet50 in transfer learning, this study has chosen to further optimize and enhance it. However, we face specific challenges when dealing with imbalanced breast cancer datasets. Firstly, the imbalance in the dataset tends to bias the model towards categories with a larger number of instances, potentially leading to a significant reduction in the detection rate of key but less frequent categories, especially in early-stage breast cancer cases, thereby increasing the risk of missed diagnoses. Secondly, given the limited sample size of minority categories, there is a risk of overfitting in these samples, affecting the model’s generalization ability for novel and previously unseen minority class samples. Therefore, the goal of this study is to overcome these challenges by meticulously optimizing the application of ResNet50 in transfer learning, thereby enhancing the model’s overall performance in dealing with imbalanced breast cancer data.
2.3 Attention mechanism
In the field of medical imaging, the attention mechanism [32] in deep learning has proven to be a groundbreaking technique. It enables models to assign varying levels of importance to different regions within medical images, such as MRI scans, X-rays, CT scans, and histopathology images [33]. This targeted focus allows the models to prioritize and analyze the most pertinent areas for diagnosing diseases, identifying anomalies, or tracking the progression of medical conditions. By concentrating on crucial features within these images, the attention mechanism greatly enhances the accuracy and efficiency of automated medical image analysis, aiding in early detection and improved patient outcomes. In the context of breast cancer image classification, introducing the attention mechanism approach can prompt the model to learn critical and discriminative features in breast cancer images that are crucial for accurately distinguishing between malignant and benign tissues. This technique can improve the model’s performance while reducing the reliance on global image feature extraction and improving the accuracy of identifying critical breast cancer regions [34].
Researchers have proposed various breast cancer segmentation and classification systems incorporating attention mechanisms. For instance, Vardhan et al. [35] developed a breast cancer segmentation system that uses convolutional neural networks with interpretable artificial intelligence and the attention mechanism to improve breast cancer detection and classification accuracy and speed. Xu et al. [36] designed a system for breast cancer segmentation that uses a modified DenseNet architecture with an attention mechanism to optimize the convolutional neural network on the BreakHis dataset for high-precision classification of breast cancer. The method was tested on the BreakHis dataset, and the method’s performance was evaluated using accuracy metrics for both binary and multi-classification. Wang et al. [37] proposed a breast cancer classification model that combines deep migration learning and visual attention mechanisms to improve the accuracy and efficiency of real-world clinical diagnosis. This deep-migration-based attention mechanisms approach uses migration learning to overcome the overfitting problem and combines the attention module to focus on characterizing breast lesions. Wang et al.’s work [37] achieves high-precision results that outperform other state-of-the-art breast image classification algorithms. This section summarizes recent scholarly research and exploration in breast cancer classification, focusing on the use of the attention mechanism. The insights gleaned indicate that applying the attention mechanism in breast cancer image classification methods can enhance the accuracy of crucial region identification and boost the interpretability of the model simultaneously. Still, the difficulty of training and parameter tuning is also increased accordingly. In this study, a lightweight design of the attention module is proposed, which incorporates adaptive average pooling, a two-layer fully connected network, and the dual activation function of LeakyReLU and ReLU to efficiently extract and enhance the critical information in the image through item-by-item multiplication (element-wise multiplication of individual pixels and features within the images)of features images, aiming to improve performance and reduce computational consumption.
3 Methods
3.1 Transfer ResNet50
The ResNet50 [9] model has a relatively complex structure, including a 50-layer deep network structure. The specific structure is shown in Table 1. The primary module is made up of numerous residual blocks. Each block is equipped with multiple convolutional layers and distinctive identity maps. Uniquely, the ResNet50 model introduces the concept of residual learning. This innovation solves the problem of vanishing gradients in deep networks, thereby opening up the possibility of building deeper neural networks than ever before. This design breakthrough allowed for achieving exceptional performance. However, considering the efficient utilization of data and computing resources, the transfer learning strategy is adopted in this paper based on the pre-trained ResNet50 model for model building. Through transfer learning, full benefit can be obtained from the high-level feature representations learned by ResNet50 on large-scale datasets in medical image classification tasks. This allows for swiftly attaining superior classification outcomes, even with constrained data and computing capacities.
Table 1: The ResNet50 Model
Module Name Output Size 50 Layers
Conv1 $112\times 112112 × 112$ $7\times 77 × 7$, 64 stride=2
[…]
Pooling $1\times 11 × 1$ Average Pool, 1000-d FC.
Note: The term "1000-d FC" in the pooling layer refers to a fully-connected layer consisting of 1000 neurons.
3.2 Dual-activated lightweight attention ResNet50
The paper proposes a dual-activation lightweight attention ResNet50 model method. Initially, a pre-trained ResNet50 is utilized as the foundational architecture. Building upon this, an innovative dual-activation lightweight attention module is introduced, along with a newly designed fully connected layer tailored for specific tasks. This attention module is integrated into the fourth residual block of ResNet50, characterized by the amalgamation of adaptive average pooling and a two-layer fully connected network, employing LeakyReLU[10] and ReLU [11] as its dual activation functions. Additionally, the newly added fully connected layer is positioned at the end of the modified ResNet50 model, primarily aimed at addressing the complexities of binary classification problems.
In the dual-activation lightweight attention module proposed in this paper, adopt the LeakyReLU activation function [10] for the initial fully connected layer, leveraging its competence in allowing for a slight gradient flow under zero. This is done to mitigate the issue of vanishing gradients, thereby enhancing the training efficacy of deep neural networks. The ReLU activation function [11] is chosen for the second fully connected network to further improve the model performance and image feature representation.
Combining these two activation functions greatly enhances the model’s expressive power and capacity to learn intricate image features, making the model more suitable for image classification tasks. In the experiments, this design improves accuracy and performance while maintaining the convergence speed of the model. The flowchart and visualization of the "Dual Activation Lightweight Attention ResNet50 Module" can be comprehended through the displayed in Fig. 2 and Fig. 3, respectively, and the overall technical framework is shown in Fig. 4.
Fig. 2: Flowchart for dual-activated lightweight attention ResNet50 model
Fig. 3: Visualization of dual-activated lightweight attention ResNet50 model
Fig. 4: The framework of dual-activated lightweight attention ResNet50 model
4 Experiment
4.1 BreakHis Dataset
The BreakHis dataset, utilized in this study focusing on breast cancer histopathological images, was generously provided by scholar Spanha [38]. This dataset comprises histopathology samples of breast cancer tumors obtained from 82 patients. The samples were sectioned under a microscope and then imaged at different magnifications (40×, 100×, 200×, 400×) to capture intricate details. Experienced medical professionals meticulously interpreted and identified the images. The dataset consists of a total of 7,909 breast cancer pathological images, comprising 2,429 benign tumor images and 5,429 malignant tumor images. Each image has dimensions of 700 × 460 pixels and is stored in PNG format with three color channels. The detailed statistics of the dataset can be found in Table 2, covering sample size, class distribution, etc. This multi-angle data resource provides a solid foundation for effective image classification and analysis research. In addition, to more intuitively display the breast cancer tissue details under different magnifications, Fig. 5 presents these pathological images as a practical example.
Table 2: BreakHis dataset: benign and malignant image distribution across magnification factors
Class Magnification Factor
40X 100X 200X 400X
Benign
625 644 623 588
Malignant
1370 1437 1390 1232
Total No of Images
1995 2081 2013 1820
Refer to caption 40x
Refer to caption 100x
Refer to caption 200x
Refer to caption 400x
Fig. 5: Images of breast cancer pathology at four magniations
4.2 Experimental environment
The experiments were carried out on an NVIDIA RTX A4000 GPU, utilizing the PyTorch 2.0.1 framework paired with Python 3.10.
4.3 Experiment implementation
This research utilizes a ratio of 6:1:3 to divide the processed dataset into training, verification, and test sets, as presented in Table 3. The training set is utilized for model training, during which parameter learning occurs. The verification set is employed to assess the model’s generalization ability. Through continuous verification and parameter fine-tuning, superior model performance is ensured, with the optimal model being retained during the verification process. The test set measures the model’s recognition rate and generalization performance. Before model training, the training data is shuffled to enhance the model’s robustness and generalization capabilities.
In deep learning, the configuration of various parameters is paramount. Particularly, the learning rate plays a crucial role; an excessively high learning rate may result in the loss value exploding, while a meager learning rate may lead to overfitting and reduce the model’s convergence speed. In previous studies, the initial learning rate is usually set in the range of 0.1 to 0.0001. Considering the relatively small size of the BreakHis medical imaging dataset (containing 7909 images) and the uneven distribution of the data, a lower learning rate of 0.0001 was chosen for this study so that the model could adjust the weights more accurately. This is important for dealing with small and unbalanced datasets to ensure that all categories are given proper attention. In addition, the setting of the training batch size and the number of iterations is extremely important. In addition, the setting of the training batch size and the number of iterations is extremely important. Previous studies usually set the batch size between 32 and 128. Considering the GPU’s 16GB video memory limit, the lowest batch size of 32 was chosen for this study. As for the number of iterations, it is also usually set between 50 and 128. Given the small size of the BreakHis dataset, this study limited the number of iterations to 50, aiming to achieve a balance between training effectiveness and computational efficiency. Finally, this study also specifically considered the setting of the dropout rate. Previous studies generally set the dropout rate between 0.25 and 0.5. In the processing of the unbalanced BreakHis dataset, a lower dropout rate helps to ensure that the categories with smaller numbers are sufficiently learned, thus improving the model’s recognition accuracy and generalization ability for these categories, thus this study chose a lower dropout rate of 0.25.
The method proposed in this study as well as the benchmark model illustrated in Table 5 apply the above settings to ensure the fairness of the experiments and the comparability of the results. These carefully selected parameters provide a solid foundation for the experimental design and ensure the reliability and efficiency of the model throughout the study, the parameter settings are shown in Table 4.
Table 3: Dataset split for experimental task
Split Percentage Purpose
Training 60% Model Training
Validation 10% Hyperparameter Tuning
Testing 30% Model Evaluation
Table 4: Experimental configuration
Parameter Value
Batch Size 32
Epochs 50
Learning Rate 0.0001
Dropout 0.25
4.4 Baseline selection
Table 5 presents a comparative analysis of various deep learning methods, encompassing both widely adopted approaches and the novel strategy introduced in this paper. The following section briefly outlines several of the models included in the comparison:
SEResNet50 [39]: Hu et al. proposed this model in 2018. The key feature of this model is integrating the "squeeze and excitation" attention mechanism, which aims to improve the network’s capabilities.
DenseNet-121 [40]: Huang et al. proposed this model in 2017 by introducing dense connections (dense blocks) to transfer and reuse information between network layers, thus improving the efficiency of feature transmission and reproduction. This study employs transfer learning with DenseNet-121 is employed for breast cancer classification. The advantage of using a pre-trained DenseNet is that it can adapt to small sample data.
VGG16 [41]: This model is from Simonyan and Zisserman and was released in 2014. It is known for its clean structure and thoughtful design philosophy and is widely used for image classification and feature extraction tasks, relying mainly on deeply stacked convolutional layers.
VGG16Inception [42]: The model is presented by M. Saini and S. Susan. The article delves into a deep learning model called "VGGIN-Net," mainly applied to analyze unbalanced breast cancer datasets. The study introduces a new deep learning framework that incorporates transfer learning to address the problem of category imbalance in biomedical datasets, applied explicitly to the BreakHis breast cancer dataset. To reduce training time and computational cost, the model uses a pre-trained network to migrate learned weights from one domain to another."
ViT (Vision Transformer) [43]: proposed by Dosovitskiy et al. in 2020, it has gained widespread attention by successfully applying the Transformer architecture to image recognition tasks. This model has shown great strength in image recognition.
Swin-Transformer [44]: Liu et al. created this model in 2021, and its performance in various vision tasks is outstanding. This model is unique in that it solves the computational and memory problems that can be encountered when processing large images.
Dinov2_Vitb14 [45]: proposed in 2023 by a team of researchers at Facebook, these models contain a pre-training dataset of 142 million unlabelled images to obtain efficient visual properties demonstrating general solid performance.
Each of the above models has uniqueness and excellent image classification and vision tasks. To emphasize the effectiveness of the proposed approach, an in-depth comparison of all these models has been conducted, achieving significant improvements in various evaluation metrics.
Table 5: Classification results of 9 network models
Network Models 40X 100X
Accuracy F1 IBA GMean Accuracy F1 IBA GMean
SEResNet50 [39]
0.883 0.779/0.883 0.860 0.860 0.850 0.720/0.849 0.813 0.812
DensNet121 [40]
0.801 0.639/0.800 0.763 0.762 0.811 0.654/0.809 0.755 0.753
VGG16 [41]
0.902 0.813/0.902 0.877 0.877 0.850 0.724/0.851 0.830 0.829
VGG16Inception [42]
0.977 0.955/0.977 0.965 0.965 0.971 0.943/0.971 0.964 0.964
ViT [43]
0.835 0.693/0.833 0.769 0.767 0.834 0.694 /0.833 0.788 0.787
Swin-Transformer [44]
0.843 0.711/0.844 0.819 0.818 0.877 0.767/0.876 0.836 0.835
Dinov2_Vitb14 [45]
0.873 0.762/0.873 0.843 0.843 0.869 0.753/0.868 0.835 0.834
RseNet50 [9]
0.927 0.860/0.927 0.914 0.914 0.944 0.891/0.944 0.922 0.922
Dual-Activated Lightweight Attention ResNet50 Model 0.985 0.970/0.985 0.980 0.980 0.987 0.974/0.987 0.985 0.985
Network Models 200X 400X
Accuracy F1 IBA GMean Accuracy F1 IBA GMean
SEResNet50 [39]
0.906 0.820/0.906 0.876 0.875 0.876 0.770/0.878 0.864 0.863
DensNet121 [40]
0.812 0.659/0.812 0.775 0.774 0.779 0.603/0.777 0.737 0.736
VGG16 [41]
0.851 0.722/0.850 0.812 0.811 0.848 0.718/0.848 0.817 0.817
VGG16Inception [42]
0.954 0.910/0.954 0.938 0.938 0.889 0.788/0.888 0.857 0.856
ViT [43]
0.841 0.705/0.840 0.801 0.800 0.832 0.691/0.831 0.803 0.803
Swin-Transformer [44]
0.876 0.766/0.875 0.838 0.837 0.846 0.724/0.851 0.791 0.789
Dinov2_Vitb14 [45]
0.853 0.726/0.852 0.802 0.800 0.885 0.782/0.885 0.840 0.839
RseNet50 [9]
0.960 0.922/0.960 0.947 0.947 0.925 0.856/0.925 0.918 0.918
Dual-Activated Lightweight Attention ResNet50 Model 0.979 0.958/0.979 0.98 0.98 0.943 0.891/0.944 0.945 0.945
Note: In the table, the DenseNet-121 model indicates that employed a pre-trained DenseNet-121 model for transfer learning.
Table 6: The count of parameters in the network model and the time it takes to converge.
Network Models Convergence Time (seconds) Total Parameters Average Time
40X 100X 200X 400X
SEResNet50 [39]
2696.22 2814.30 2711.89 2356.18 27073090 2644.65
DensNet121 [40]
2352.46 2643.79 2423.36 2078.34 9028970 2347.49
VGG16 [41]
3212.95 3182.54 3011.53 2640.19 134268738 3011.80
VGG16Inception [42]
3074.99 3253.50 3303.12 2682.61 38139938 3078.56
ViT [43]
2932.11 3022.54 3013.86 2691.16 51375106 2914.92
Swin-Transformer [44]
3020.11 3132.42 2940.98 2592.96 86745274 2921.62
Dinov2_Vitb14 [45]
2895.02 2303.25 2245.98 2017.15 86582018 2365.35
RseNet50 [9]
2920.53 2933.26 2307.94 2097.33 21797672 2564.77
Dual-Activated Lightweight Attention ResNet50 Model 2133.02 2319.50 1990.58 1896.82 25611842 2084.98
Note: In the table, the DenseNet-121 model indicates that employed a pre-trained DenseNet-121 model for transfer learning.
4.5 Experiment metrics
Used four metrics, Accuracy [46, 47, 48], F1 Score [49], IBA [50], and GMean [51], to assess the performance of the model holistically. The F1 Score is a crucial metric that combines accuracy and recall, providing a comprehensive measure of a model’s ability to classify positive and negative instances. This is especially useful for imbalanced datasets where positive and negative samples differ significantly. [52]. IBA (Imbalance Balanced Accuracy) takes a targeted approach to tackle the class imbalance challenge by harmonizing accuracy and balance. This approach offers a more holistic assessment of the model’s performance, encompassing sample diversity and fairness [53]. GMean (Geometric Mean) is a metric that calculates the geometric mean of individual classification accuracies, which is used to evaluate the model’s ability to handle category imbalance and ensure that the model can perform well across all classes [54].
In Table 5 showcases the performance of the model on the test set across various image resolutions, such as 40X, 100X, 200X, and 400X. Regardless of the solution, the model consistently outperforms nine other popular network models regarding accuracy, F1 score, IBA, and GMean. These results demonstrate that the network model has a significant advantage regardless of which image resolution is used. This gives us a deep insight into the model’s excellent performance and great potential in performing medical image classification tasks.
In Table 6, presents a comparison of the model with eight other network models regarding the number of model parameters and convergence speed. Notably, the DenseNet-121 model possesses the fewest parameters. However, the model has the shortest convergence time and the fastest average convergence speed across all image resolutions, including 40X, 100X, 200X, and 400X. This further reinforces the superior performance and potential of the model for medical image classification tasks. These results proved the efficiency and effectiveness of the model.
5 Performance evaluation and visualization
To thoroughly oversee the model’s performance and glean insights into the training process, focus on several key metrics, including training loss, validation loss, training accuracy, and validation accuracy. Visualizing these metrics can demonstrate the model’s training performance. The results of analyzing these key metrics are presented in detail below.
Refer to caption 40X
Refer to caption 100X
Refer to caption 200X
Refer to caption 400X
Fig. 6: Training and validation loss with error bars
5.1 Loss analysis
This study utilized a dual-activated lightweight attention ResNet50 model for breast cancer classification. The critical metrics for monitoring model performance and understanding the training process are training loss and validation loss [55, 56, 57]. These loss values play an important role when evaluating model performance, and their visualization gives us a clearer picture of the training progress.
Fig. 6 shows the variation of training loss and validation loss with the number of training iterations for different image resolutions (including 40X, 100X, 200X, and 400X). The x-axis denotes the quantity of training iterations, while the y-axis signifies the values of the loss function. Used line plots to show the trend of training loss and validation loss and added error bars to each data point to show the standard deviation of the loss values.
As training progresses, a gradual decrease in training loss and validation loss is observed, which signifies that the model is effectively learning and adapting to the data. This is a positive trend and shows that the model is making progress on the task. In the later stages of training, the validation loss changes less. It offers a stable trend, indicating that the model steers from overfitting on the validation set and demonstrates strong generalization to new data.
Refer to caption 40X
Refer to caption 100X
Refer to caption 200X
Refer to caption 400X
Fig. 7: Training and validation accuracy with error bars
5.2 Accuracy analysis
In Fig. 7, the evolution of training accuracy and validation accuracy at different image resolutions (40X, 100X, 200X, and 400X) is depicted [58]. The horizontal axis represents the number of training iterations, while the vertical axis represents the accuracy. A line graph is employed to portray the accuracy trend, and error bars are incorporated to convey the standard deviation of the accuracy values.
After meticulous analysis, the following outcomes have been derived: Training and validation accuracy undergo a gradual ascent, showcasing the model’s persistent learning and adaptation throughout the training process. This upward trajectory suggests that the model is advancing in comprehending and classifying the data. Additionally, the validation accuracy maintains stability, especially in the later stages of training, implying that the model avoids overfitting to the training data and can generalize effectively to unseen data. These findings are pivotal in comprehending the model’s performance and can facilitate further optimization for enhanced breast cancer classification results.
6 Summary and conclusions
This paper focuses on applying a lightweight attention mechanism to a ResNet50 model for classifying medical histopathology tissue images, mainly when dealing with limited-scale datasets. The ResNet50 model incorporates the lightweight attention mechanism and a dropout layer to optimize the pre-trained deep network layer. The performance evaluation includes metrics such as accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of this paper model’s performance compared to existing techniques. The experimental results highlight significant advantages, particularly when addressing the challenges of unbalanced breast cancer datasets. The model demonstrates improved robustness, generalization, faster convergence during training, and effective prevention of overfitting. These findings provide a solid basis for enhancing classification accuracy. In future research, investigating more breast multiclassification problems is planned.
[…]

### v6

@@
-1] organization=School of Computer Sciences , addressline=Universiti Sains Malaysia, city=11800 USM Penang, country=Malaysia
+
 Abstract
-Despite the remarkable results of deep learning in breast cancer histopathology image classification, challenges such as data imbalance and interpretability still exist and require cross-domain knowledge and collaboration among medical experts.
-This study proposes a breast cancer classification method using a dual-activated lightweight attention ResNet50 model, effectively addressing data imbalance and interpretability challenges.
-The model fuses a pre-trained deep ResNet50 and a lightweight attention mechanism to accomplish classification by embedding an attention module in layer 4 of ResNet50 and adding two fully connected layers.
-The fully connected network design employs LeakyReLU and ReLU activation functions.
-The model outperforms SEResNet50, DensNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50 models regarding precision, accuracy, recall, F1 score, and GMean, especially in the application performance on the BreakHis dataset.
-In particular, the model demonstrates significant robustness and broad applicability when dealing with the unbalanced breast cancer dataset.
-The model has been evaluated on histopathology images at magnification factors of 40X, 100X, 200X, and 400X, achieving accuracies of 98.5%, 98.7%, 97.9%, and 94.3%, respectively.
-The study comprehensively assessed the model’s performance.
-In the later stages of training, the validated losses and accuracies change minimally, showing that the model avoids overfitting and exhibits good generalization ability.
-This model exhibited the fastest convergence in all laboratory experiments, even though its parameters are not the smallest.
-This highlights the model’s efficacy as a lightweight attention framework, showcasing its efficiency in achieving rapid convergence without compromising performance.
-keywords:
+Automatic breast cancer classification technology in histopathology images plays a crucial role in accurate diagnosis and treatment planning.
+Recently, histopathology image classification approaches based on ResNet architecture have been widely proposed as they vastly improve the classification accuracy using skip connection to prevent vanishing gradient issues by combining low-level and high-level feature information.
+However, traditional ResNet architecture suffers from challenges such as data imbalance and interpretability, which still require cross-domain knowledge and collaboration among medical experts.
+This study proposes a new breast cancer classification method using a Dual-Activated Lightweight Attention ResNet50 (DALAResNet50) approach to address data imbalance and interpretability challenges effectively.
+The proposed approach combines a pre-trained deep ResNet50 method and a lightweight attention mechanism to accomplish the classification by embedding an attention module in layer 4 of ResNet50 and adding two fully connected layers based on LeakyReLU and ReLU activation functions to enhance the ability of feature learning.
+The proposed DALAResNet50 approach has been evaluated on breast cancer histopathology images derived from BreakHis Datase at magnification factors of 40X, 100X, 200X, and 400X, achieving accuracies of 98.5%, 98.7%, 97.9%, and 94.3%, respectively.
+The DALAResNet50 approach was comprehensively assessed with popular deep learning approaches such as SEResNet50, DensNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50 models.
+The reported results of DALAResNet50 outperform the compared approaches regarding precision, accuracy, recall, F1 score, and GMean, demonstrating significant robustness and broad applicability when dealing with different magnifications and unbalanced breast cancer datasets.
+Index Terms:
+Lightweight attention mechanism, Convolutional Neural Networks, breast cancer classification.
 […]
-1 Introduction
+=-21pt
+Figure 1: Overview of Convolutional Neural Network
+I Introduction
 Breast cancer, particularly among women, has emerged as one of the leading causes of fatal diseases, with its incidence on a continuous rise, eliciting worldwide health concerns [1, 2].
@@
 This profound learning advancement has revolutionized breast cancer diagnosis [4, 5, 6, 7].
-One key innovation in this area is developing and applying Convolutional Neural Networks (CNNs), a specific class of deep learning models designed to process data with a grid-like structure, such as medical images.
+One key innovation in this area is developing and applying Convolutional Neural Networks (CNNs), a specific class of deep learning models designed to process data, such as medical images, with a grid-like structure.
 The concept of CNN was first introduced back in 1998 [8].
 The concept describes a method for efficiently classifying image data.
-CNNs gradually extract abstract features of an image through multi-level convolutional and pooling operations, allowing the network to extract necessary features from different categories of images so that the different categories can be classified accurately based on those features.
+CNNs gradually extract abstract features of an image through multi-level convolutional and pooling operations, allowing the network to remove necessary features from different categories of pictures so that the other categories can be classified accurately based on those features.
 CNNs have three main layers: convolutional, pooling, and fully connected.
@@
 1.
-However, despite the significant achievements of deep learning in medical image classification, conventional CNN architectures struggle with unbalanced medical image data.
+However, despite the considerable achievements of deep learning in medical image classification, conventional CNN architectures struggle with unbalanced medical image data.
 To tackle this issue, In this paper, a lightweight attention mechanism based on ResNet50 [9] is introduced to address the challenge of unbalanced medical image data in conventional CNN architectures.
@@
 (b) Integrating the attention mechanism module and ResNet50 to form a feature-rich model, the model achieves more efficient feature extraction based on fusing the two techniques.
-Specifically, in the fully connected network part of the attention mechanism module, we employed LeakyReLU [10] and ReLU [11] activation functions.
-Leaky-ReLU was chosen to mitigate the gradient vanishing problem, while ReLU was used to enhance the image feature representation, thus better capturing multi-scale spatial patterns.
+Specifically, in the fully connected network component of the attention mechanism module, LeakyReLU [10] and ReLU [11] activation functions were utilized.
+LeakyReLU was selected for its ability to mitigate the gradient vanishing problem.
+In contrast, ReLU was employed to enhance the representation of image features.
+This choice was instrumental in better capturing multi-scale spatial patterns.
 (c) In this paper, test images with different magnifications and verify that the network is invariant to image scales and performs better than other state-of-the-art methods.
@@
 Finally, Section 6 provides a summary and conclusions.
-Fig. 1: Overview of Convolutional Neural Network
-2 Related works
+II Related works
 Currently, two main methods are primarily used for cancer tissue histopathology image classification: traditional manual feature extraction and deep learning-based methods.
@@
 The following section delves into these techniques in detail.
-2.1 CNN
+II-A CNN
 CNN [8], one of the most prevalent deep learning techniques, is apt for tasks involving image processing [14].
@@
 This surpassed traditional methods, though parameter optimization and validation on unbalanced datasets require further exploration.
-The Inception model is a neural network architecture introduced by Google, which aims to solve the computation problem and several parameters in neural networks and is suitable for image classification and related tasks.
+The Inception model is a neural network architecture introduced by Google that aims to solve the computation problem and several parameters in neural networks and is suitable for image classification and related tasks.
 The hybrid model based on CNN-Inception-V4 was used by Nazir et al.
@@
 Given the significant superiority of the ResNet50 model over other methods in terms of accuracy, this paper has decided to modify the ResNet50 model to address these specific challenges.
-2.2 Transfer learning
-Transfer learning applies pre-trained models from one image domain to another, effectively bridging the gap between different image domains.
-In the context of breast cancer image classification, this pre-trained model approach involves fine-tuning CNN models, which were initially trained on a comprehensive large-scale image data set, and then fine-tuning it to adapt to the data set specifically used for breast cancer detection.
-Through this step, the model is optimized to better meet the specific needs of breast cancer classification.
+II-B Transfer learning
+Transfer learning applies pre-trained models from one image domain to another, bridging the gap between different image domains.
+In the context of breast cancer image classification, this pre-trained model approach involves fine-tuning CNN models, which were initially trained on a comprehensive large-scale image data set, and then fine-tuning it to adapt to the data set used explicitly for breast cancer detection.
+This step optimizes the model better to meet the specific needs of breast cancer classification.
 By using the inherent ability of the pre-trained model in generalization and feature extraction, this strategy significantly improves the efficiency of training and the accuracy of classification performance [23, 24, 25, 26, 27, 28, 29].
@@
 After training and testing these models on a dataset of 2453 breast histopathology images, they found that ResNet50 had a superior performance with an accuracy rate of 90.2%, an Area Under the Curve (AUC) of 90.0%, and a recall rate of 94.7%.
-In conclusion, considering the exceptional robustness of ResNet50 in transfer learning, this study has chosen to further optimize and enhance it.
-However, we face specific challenges when dealing with imbalanced breast cancer datasets.
+In conclusion, this study has focused on further optimization and enhancement due to the exceptional robustness of ResNet50 in transfer learning.
+Nevertheless, the investigation encounters specific challenges when dealing with imbalanced breast cancer datasets.
 Firstly, the imbalance in the dataset tends to bias the model towards categories with a larger number of instances, potentially leading to a significant reduction in the detection rate of key but less frequent categories, especially in early-stage breast cancer cases, thereby increasing the risk of missed diagnoses.
 Secondly, given the limited sample size of minority categories, there is a risk of overfitting in these samples, affecting the model’s generalization ability for novel and previously unseen minority class samples.
-Therefore, the goal of this study is to overcome these challenges by meticulously optimizing the application of ResNet50 in transfer learning, thereby enhancing the model’s overall performance in dealing with imbalanced breast cancer data.
-2.3 Attention mechanism
-In the field of medical imaging, the attention mechanism [32] in deep learning has proven to be a groundbreaking technique.
+Therefore, this study aims to overcome these challenges by meticulously optimizing the application of ResNet50 in transfer learning, thereby enhancing the model’s overall performance in dealing with imbalanced breast cancer data.
+II-C Attention mechanism
+In medical imaging, the attention mechanism [32] in deep learning has proven to be a groundbreaking technique.
 It enables models to assign varying levels of importance to different regions within medical images, such as MRI scans, X-rays, CT scans, and histopathology images [33].
 This targeted focus allows the models to prioritize and analyze the most pertinent areas for diagnosing diseases, identifying anomalies, or tracking the progression of medical conditions.
-By concentrating on crucial features within these images, the attention mechanism greatly enhances the accuracy and efficiency of automated medical image analysis, aiding in early detection and improved patient outcomes.
-In the context of breast cancer image classification, introducing the attention mechanism approach can prompt the model to learn critical and discriminative features in breast cancer images that are crucial for accurately distinguishing between malignant and benign tissues.
+By concentrating on crucial features within these images, the attention mechanism dramatically enhances the accuracy and efficiency of automated medical image analysis, aiding in early detection and improved patient outcomes.
+In the context of breast cancer image classification, introducing the attention mechanism approach can prompt the model to learn critical and discriminative features in breast cancer images crucial for accurately distinguishing between malignant and benign tissues.
 This technique can improve the model’s performance while reducing the reliance on global image feature extraction and improving the accuracy of identifying critical breast cancer regions [34].
@@
 This section summarizes recent scholarly research and exploration in breast cancer classification, focusing on the use of the attention mechanism.
-The insights gleaned indicate that applying the attention mechanism in breast cancer image classification methods can enhance the accuracy of crucial region identification and boost the interpretability of the model simultaneously.
-Still, the difficulty of training and parameter tuning is also increased accordingly.
-In this study, a lightweight design of the attention module is proposed, which incorporates adaptive average pooling, a two-layer fully connected network, and the dual activation function of LeakyReLU and ReLU to efficiently extract and enhance the critical information in the image through item-by-item multiplication (element-wise multiplication of individual pixels and features within the images)of features images, aiming to improve performance and reduce computational consumption.
-3 Methods
-3.1 Transfer ResNet50
+The insights gleaned indicate that applying the attention mechanism in breast cancer images and classification methods can simultaneously enhance the accuracy of crucial region identification and boost the interpretability of the model.
+However, this also correspondingly increases the following difficulties:
+II-C1 Limitations in global feature extraction
+Although the attention mechanism is designed to enhance model performance by focusing on critical parts of the image, thereby playing a positive role in strengthening the model’s ability to process global information, it may still face challenges in some implementations or specific application scenarios, unable to fully capture the international contextual details of the entire image.
+II-C2 Insufficiency in feature priority precision
+Lacking a sophisticated method to discern the significance of diverse features, attention mechanisms might uniformly weigh all channel information.
+This approach risks overlooking vital features essential for precise classification, undermining the model’s ability to highlight and utilize discriminative patterns within the data effectively.
+II-C3 Inefficiency of Activation Functions
+The reliance on ReLU within attention mechanisms can lead to the ”dying ReLU” issue, wherein neurons fail to activate over specific data ranges, stalling the learning progression.
+Such constraints significantly diminish the model’s adaptability during the training phase, adversely affecting the dynamics and overall efficacy of the attention-enhanced learning process.
+II-C4 Lack of feature adaptability
+Attention mechanisms may exhibit rigidity in adjusting and enhancing features non-linearly according to the unique characteristics of each input.
+This inflexibility can constrain the model’s ability to adapt and optimally leverage the diverse information within the data, thereby reducing the overall efficiency in feature utilization.
+This study designed an innovative lightweight attention module to overcome the identified difficulties.
+This design incorporates adaptive average pooling, a two-layer fully connected network for channel recalibration, dual activation functions of LeakyReLU [10] and ReLU [11], and element-wise multiplication.
+Adaptive average pooling tackles the limitations in global feature extraction, whereas the two-layer fully connected network for channel recalibration optimizes feature priority precision.
+The combination of LeakyReLU and ReLU effectively addresses the inefficiencies of a single activation function, enhancing activation function efficiency, and element-wise multiplication improves feature adaptability.
+These pioneering approaches synergistically aim to significantly elevate model performance while mitigating computational resource requirements, providing a more efficient and adaptable solution for integrating attention mechanisms into breast cancer image classification models.
+III Methods
+III-A Transfer ResNet50
 The ResNet50 [9] model has a relatively complex structure, including a 50-layer deep network structure.
-The specific structure is shown in Table 1.
+The specific structure is shown in Table I.
 The primary module is made up of numerous residual blocks.
@@
 This allows for swiftly attaining superior classification outcomes, even with constrained data and computing capacities.
-Table 1: The ResNet50 Model
+TABLE I: The ResNet50 Model
 Module Name Output Size 50 Layers
@@
 Pooling $1\times 11 × 1$ Average Pool, 1000-d FC.
-Note: The term "1000-d FC" in the pooling layer refers to a fully-connected layer consisting of 1000 neurons.
-3.2 Dual-activated lightweight attention ResNet50
-The paper proposes a dual-activation lightweight attention ResNet50 model method.
-Initially, a pre-trained ResNet50 is utilized as the foundational architecture.
-Building upon this, an innovative dual-activation lightweight attention module is introduced, along with a newly designed fully connected layer tailored for specific tasks.
-This attention module is integrated into the fourth residual block of ResNet50, characterized by the amalgamation of adaptive average pooling and a two-layer fully connected network, employing LeakyReLU[10] and ReLU [11] as its dual activation functions.
-Additionally, the newly added fully connected layer is positioned at the end of the modified ResNet50 model, primarily aimed at addressing the complexities of binary classification problems.
-In the dual-activation lightweight attention module proposed in this paper, adopt the LeakyReLU activation function [10] for the initial fully connected layer, leveraging its competence in allowing for a slight gradient flow under zero.
+Note: The term ”1000-d FC” in the pooling layer refers to a fully-connected layer consisting of 1000 neurons.
+TABLE II: Comparison between DALAResNet50 model and standard ResNet model
+Feature
+DALAResNet50 model
+Standard ResNet model
+Adaptive Average Pooling
+Adaptive average pooling addresses ”limitations in global feature extraction” by enhancing the model’s capacity to process global contextual information, significantly improving performance in tasks requiring comprehensive image understanding.
+Not available
+Channel Recalibration
+A two-layer fully connected network for channel recalibration optimizes ”feature priority precision,” enhancing the model’s ability to identify and focus on the most relevant features for accurate classification.
+Not available
+Activation Functions
+The dual activation functions of LeakyReLU and ReLU address the ”inefficiency of activation functions.” By incorporating LeakyReLU, the network avoids the dying ReLU problem, as it allows a small, non-zero gradient when the unit is not active, ensuring continuous learning dynamics and preventing neurons from becoming inactive across all network layers.
+This combination maintains the model’s learning capability, potentially enhancing the training process.
+ReLU only
+Element-wise multiplication
+Element-wise multiplication, as denoted by ’Scale’ in Fig.
+3, enhances ”Feature Adaptability” by allowing dynamic, non-linear adjustment of features according to a specific input, improving the network’s flexibility and utilization of information.
+Not available
+Fig. 2: Flowchart for standard ResNet model
+Fig. 3: Flowchart for DALAResNet50 model
+Fig. 4: Visualization of DALAResNet50 model
+Fig. 5: The framework of DALAResNet50 model
+III-B Dual-activated lightweight attention ResNet50
+A new approach for automatic breast cancer classification is proposed in this research using a new version of the ResNet50 technique called Dual-activated lightweight attention ResNet50 (DALAResNet50).
+The DALAResNet50 approach incorporates ResNet50, Global pooling, channel recalibration, Dual Activation functions, and the newly designed fully connected layer into a unified approach to overcome the drawbacks of the traditional ResNet50 model.
+The pre-trained ResNet50 is first utilized as a baseline architecture.
+Subsequently, the attention module is integrated into the fourth residual block of ResNet50 to pay attention to the small-scale nuclei cells of breast cancer.
+The attention mechanism improves the ability of feature learning by extracting the silent semantic information to overcome the class imbalance issue.
+The proposed attention module is characterized by the amalgamation of adaptive global pooling and a two-layer fully connected network, employing LeakyReLU [10] and ReLU [11] as dual activation functions.
+The traditional ResNet50 uses conventional pooling layers to reduce the spatial dimensions of the feature maps while retaining important information.
+However, the conventional pooling layers operate over local regions of the input feature maps, reducing their spatial dimensions but discarding spatial information, making the ResNet50 sensitive to spatial translations of the input feature map and increasing the number of parameters in the network.
+Therefore, this research proposed average-based global pooling that calculates the average value of each feature map across all spatial locations, effectively summarizing each feature’s presence throughout the entire feature map.
+The main advantage of global pooling is that it provides a fixed-size representation of the whole feature map, making the ResNet50 less sensitive to spatial translations of the input and reducing the total number of parameters in the network.
+Additionally, the DALAResNet50 adopts the LeakyReLU activation function [10] for the initial fully connected layer, leveraging its competence in allowing for a slight gradient flow under zero.
 This is done to mitigate the issue of vanishing gradients, thereby enhancing the training efficacy of deep neural networks.
-The ReLU activation function [11] is chosen for the second fully connected network to further improve the model performance and image feature representation.
-Combining these two activation functions greatly enhances the model’s expressive power and capacity to learn intricate image features, making the model more suitable for image classification tasks.
-In the experiments, this design improves accuracy and performance while maintaining the convergence speed of the model.
-The flowchart and visualization of the "Dual Activation Lightweight Attention ResNet50 Module" can be comprehended through the displayed in Fig.
+The ReLU activation function [11] is further chosen for the second fully connected network to improve the model performance and image feature representation.
+Combining these two activation functions enhances the model’s expressive power and capacity to learn intricate image features, making the model more suitable for breast cancer image classification tasks.
+Finally, the newly added fully connected layer is positioned at the end of the DALAResNet50 approach to address the complexity of binary classification problems.
+Fig.
 2 and Fig.
-3, respectively, and the overall technical framework is shown in Fig.
-4.
-Fig. 2: Flowchart for dual-activated lightweight attention ResNet50 model
-Fig. 3: Visualization of dual-activated lightweight attention ResNet50 model
-Fig. 4: The framework of dual-activated lightweight attention ResNet50 model
-4 Experiment
-4.1 BreakHis Dataset
+3 illustrate the flowcharts of the standard ResNet approach and the DALAResNet50 approach, respectively.
+These flowcharts highlight the newly added features in the DALAResNet50.
+A comprehensive comparative analysis of these differences is detailed in Table 2.
+The visualization of the DALAResNet50 approach and the entire technical framework are depicted in Fig.
+4 and Fig.
+5.
+IV Experiment
+IV-A BreakHis Dataset
 The BreakHis dataset, utilized in this study focusing on breast cancer histopathological images, was generously provided by scholar Spanha [38].
@@
 Each image has dimensions of 700 × 460 pixels and is stored in PNG format with three color channels.
-The detailed statistics of the dataset can be found in Table 2, covering sample size, class distribution, etc.
+The detailed dataset statistics in Table 3 cover sample size, class distribution, etc.
 This multi-angle data resource provides a solid foundation for effective image classification and analysis research.
@@
 6 presents these pathological images as a practical example.
-Table 2: BreakHis dataset: benign and malignant image distribution across magnification factors
+Refer to caption 40x
+Refer to caption 100x
+Refer to caption 200x
+Refer to caption 400x
+Fig. 6: Images of breast cancer pathology at four magniations
+TABLE III: BreakHis dataset: benign and malignant image distribution across magnification factors
 Class Magnification Factor
@@
 1995 2081 2013 1820
-Refer to caption 40x
-Refer to caption 100x
-Refer to caption 200x
-Refer to caption 400x
-Fig. 5: Images of breast cancer pathology at four magniations
-4.2 Experimental environment
-The experiments were carried out on an NVIDIA RTX A4000 GPU, utilizing the PyTorch 2.0.1 framework paired with Python 3.10.
-4.3 Experiment implementation
-This research utilizes a ratio of 6:1:3 to divide the processed dataset into training, verification, and test sets, as presented in Table 3.
-The training set is utilized for model training, during which parameter learning occurs.
-The verification set is employed to assess the model’s generalization ability.
-Through continuous verification and parameter fine-tuning, superior model performance is ensured, with the optimal model being retained during the verification process.
-The test set measures the model’s recognition rate and generalization performance.
-Before model training, the training data is shuffled to enhance the model’s robustness and generalization capabilities.
-In deep learning, the configuration of various parameters is paramount.
-Particularly, the learning rate plays a crucial role; an excessively high learning rate may result in the loss value exploding, while a meager learning rate may lead to overfitting and reduce the model’s convergence speed.
-In previous studies, the initial learning rate is usually set in the range of 0.1 to 0.0001.
-Considering the relatively small size of the BreakHis medical imaging dataset (containing 7909 images) and the uneven distribution of the data, a lower learning rate of 0.0001 was chosen for this study so that the model could adjust the weights more accurately.
-This is important for dealing with small and unbalanced datasets to ensure that all categories are given proper attention.
-In addition, the setting of the training batch size and the number of iterations is extremely important.
-In addition, the setting of the training batch size and the number of iterations is extremely important.
-Previous studies usually set the batch size between 32 and 128.
-Considering the GPU’s 16GB video memory limit, the lowest batch size of 32 was chosen for this study.
-As for the number of iterations, it is also usually set between 50 and 128.
-Given the small size of the BreakHis dataset, this study limited the number of iterations to 50, aiming to achieve a balance between training effectiveness and computational efficiency.
-Finally, this study also specifically considered the setting of the dropout rate.
-Previous studies generally set the dropout rate between 0.25 and 0.5.
-In the processing of the unbalanced BreakHis dataset, a lower dropout rate helps to ensure that the categories with smaller numbers are sufficiently learned, thus improving the model’s recognition accuracy and generalization ability for these categories, thus this study chose a lower dropout rate of 0.25.
-The method proposed in this study as well as the benchmark model illustrated in Table 5 apply the above settings to ensure the fairness of the experiments and the comparability of the results.
-These carefully selected parameters provide a solid foundation for the experimental design and ensure the reliability and efficiency of the model throughout the study, the parameter settings are shown in Table 4.
-Table 3: Dataset split for experimental task
+TABLE IV: Dataset split for experimental task
 Split Percentage Purpose
@@
 Testing 30% Model Evaluation
-Table 4: Experimental configuration
+TABLE V: Experimental configuration
 Parameter Value
@@
 Dropout 0.25
-4.4 Baseline selection
-Table 5 presents a comparative analysis of various deep learning methods, encompassing both widely adopted approaches and the novel strategy introduced in this paper.
-The following section briefly outlines several of the models included in the comparison:
-SEResNet50 [39]: Hu et al. proposed this model in 2018. The key feature of this model is integrating the "squeeze and excitation" attention mechanism, which aims to improve the network’s capabilities.
-DenseNet-121 [40]: Huang et al.
-proposed this model in 2017 by introducing dense connections (dense blocks) to transfer and reuse information between network layers, thus improving the efficiency of feature transmission and reproduction.
-This study employs transfer learning with DenseNet-121 is employed for breast cancer classification.
-The advantage of using a pre-trained DenseNet is that it can adapt to small sample data.
-VGG16 [41]: This model is from Simonyan and Zisserman and was released in 2014.
-It is known for its clean structure and thoughtful design philosophy and is widely used for image classification and feature extraction tasks, relying mainly on deeply stacked convolutional layers.
-VGG16Inception [42]: The model is presented by M.
-Saini and S.
-Susan.
-The article delves into a deep learning model called "VGGIN-Net," mainly applied to analyze unbalanced breast cancer datasets.
-The study introduces a new deep learning framework that incorporates transfer learning to address the problem of category imbalance in biomedical datasets, applied explicitly to the BreakHis breast cancer dataset.
-To reduce training time and computational cost, the model uses a pre-trained network to migrate learned weights from one domain to another."
-ViT (Vision Transformer) [43]: proposed by Dosovitskiy et al.
-in 2020, it has gained widespread attention by successfully applying the Transformer architecture to image recognition tasks.
-This model has shown great strength in image recognition.
-Swin-Transformer [44]: Liu et al.
-created this model in 2021, and its performance in various vision tasks is outstanding.
-This model is unique in that it solves the computational and memory problems that can be encountered when processing large images.
-Dinov2_Vitb14 [45]: proposed in 2023 by a team of researchers at Facebook, these models contain a pre-training dataset of 142 million unlabelled images to obtain efficient visual properties demonstrating general solid performance.
-Each of the above models has uniqueness and excellent image classification and vision tasks.
-To emphasize the effectiveness of the proposed approach, an in-depth comparison of all these models has been conducted, achieving significant improvements in various evaluation metrics.
-Table 5: Classification results of 9 network models
+IV-B Experimental environment
+The experiments were carried out on an NVIDIA RTX A4000 GPU, utilizing the PyTorch 2.0.1 framework paired with Python 3.10.
+IV-C Experiment implementation
+This research utilizes a ratio of 6:1:3 to divide the processed dataset into training, verification, and test sets, as presented in Table 4.
+The training set is used for model training, during which parameter learning occurs.
+The verification set is employed to assess the model’s generalization ability.
+Continuous verification and parameter fine-tuning ensure superior model performance, with the optimal model pre-trained during verification.
+The test set measures the model’s recognition rate and generalization performance.
+Before model training, the training data is shuffled to enhance the model’s robustness and generalization capabilities.
+In deep learning, the configuration of various parameters is paramount.
+Notably, the learning rate plays a crucial role; an excessively high learning rate may result in the loss value exploding, while a meager learning rate may lead to overfitting and reduce the model’s convergence speed.
+In previous studies, the initial learning rate is usually set in the range of 0.1 to 0.0001.
+Considering the relatively small size of the BreakHis medical imaging dataset (containing 7909 images) and the uneven distribution of the data, a lower learning rate of 0.0001 was chosen for this study so that the model could adjust the weights more accurately.
+This is important for dealing with small and unbalanced datasets to ensure that all categories are given proper attention.
+In addition, setting the training batch size and the number of iterations is significant.
+In addition, setting the training batch size and the number of iterations is substantial.
+Previous studies usually set the batch size between 32 and 128.
+Considering the GPU’s 16GB video memory limit, the lowest batch size of 32 was chosen for this study.
+As for the number of iterations, it is also usually set between 50 and 128.
+Given the size of the BreakHis dataset, this study limited the number of iterations to 50, aiming to balance training effectiveness and computational efficiency.
+Finally, this study also specifically considered the dropout rate setting.
+Previous studies generally set the dropout rate between 0.25 and 0.5.
+In processing the unbalanced BreakHis dataset, a lower dropout rate helps ensure that the categories with smaller numbers are sufficiently learned, thus improving the model’s recognition accuracy and generalization ability.
+Therefore, this study chose a lower dropout rate of 0.25.
+The method proposed in this study, and the benchmark model illustrated in the Table apply the above settings to ensure the experiments’ fairness and the comparability of the results.
+These carefully selected parameters provide a solid foundation for the experimental design and ensure the reliability and efficiency of the model throughout the study; the parameter settings are shown in Table 5.
+TABLE VI: Classification results of 9 network models
 Network Models 40X 100X
@@
 0.927 0.860/0.927 0.914 0.914 0.944 0.891/0.944 0.922 0.922
-Dual-Activated Lightweight Attention ResNet50 Model 0.985 0.970/0.985 0.980 0.980 0.987 0.974/0.987 0.985 0.985
+DALAResNet50 Model 0.985 0.970/0.985 0.980 0.980 0.987 0.974/0.987 0.985 0.985
 Network Models 200X 400X
@@
 0.960 0.922/0.960 0.947 0.947 0.925 0.856/0.925 0.918 0.918
-Dual-Activated Lightweight Attention ResNet50 Model 0.979 0.958/0.979 0.98 0.98 0.943 0.891/0.944 0.945 0.945
+DALAResNet50 Model 0.979 0.958/0.979 0.98 0.98 0.943 0.891/0.944 0.945 0.945
 Note: In the table, the DenseNet-121 model indicates that employed a pre-trained DenseNet-121 model for transfer learning.
-Table 6: The count of parameters in the network model and the time it takes to converge.
+TABLE VII: The count of parameters in the network model and the time it takes to converge.
 Network Models Convergence Time (seconds) Total Parameters Average Time
@@
 2696.22 2814.30 2711.89 2356.18 27073090 2644.65
-DensNet121 [40]
+DenseNet121 [40]
 2352.46 2643.79 2423.36 2078.34 9028970 2347.49
@@
 2895.02 2303.25 2245.98 2017.15 86582018 2365.35
-RseNet50 [9]
+ResNet50 [9]
 2920.53 2933.26 2307.94 2097.33 21797672 2564.77
-Dual-Activated Lightweight Attention ResNet50 Model 2133.02 2319.50 1990.58 1896.82 25611842 2084.98
-Note: In the table, the DenseNet-121 model indicates that employed a pre-trained DenseNet-121 model for transfer learning.
-4.5 Experiment metrics
+DALAResNet50 Model 2133.02 2319.50 1990.58 1896.82 25611842 2084.98
+Note: In the table, the DenseNet-121 model indicates that a pre-trained DenseNet-121 model was employed for transfer learning.
+IV-D Baseline selection
+Table 6 presents a comparative analysis of various deep learning methods, encompassing widely adopted approaches and the novel strategy introduced in this paper.
+The following section briefly outlines several of the models included in the comparison:
+SEResNet50 [39]: Hu et al. proposed this model in 2018. The key feature of this model is integrating the ”squeeze and excitation” attention mechanism, which aims to improve the network’s capabilities.
+DenseNet-121 [40]: Huang et al.
+proposed this model in 2017 by introducing dense connections (dense blocks) to transfer and reuse information between network layers, thus improving the efficiency of feature transmission and reproduction.
+This study employs transfer learning with DenseNet-121, which is utilized for classifying breast cancer.
+The advantage of using a pre-trained DenseNet is that it can adapt to small sample data.
+VGG16 [41]: This model is from Simonyan and Zisserman and was released in 2014.
+It is known for its clean structure and thoughtful design philosophy and is widely used for image classification and feature extraction tasks, relying mainly on deeply stacked convolutional layers.
+VGG16Inception [42]: The model is presented by M.
+Saini and S.
+Susan.
+The article delves into a deep learning model called ”VGGIN-Net,” which is mainly applied to analyze unbalanced breast cancer datasets.
+The study introduces a new deep learning framework that incorporates transfer learning to address the problem of category imbalance in biomedical datasets, applied explicitly to the BreakHis breast cancer dataset.
+To reduce training time and computational cost, the model uses a pre-trained network to migrate learned weights from one domain to another.”
+ViT (Vision Transformer) [43]: Proposed by Dosovitskiy et al.
+in 2020, it has gained widespread attention by successfully applying the Transformer architecture to image recognition tasks.
+This model has shown great strength in image recognition.
+Swin-Transformer [44]: Liu et al.
+created this model in 2021, and its performance in various vision tasks is outstanding.
+This model is unique in that it solves the computational and memory problems that can be encountered when processing large images.
+Dinov2_Vitb14 [45]: Proposed in 2023 by a team of researchers at Facebook, these models contain a pre-training dataset of 142 million unlabelled images to obtain efficient visual properties demonstrating general solid performance.
+Each of the above models has uniqueness and excellent image classification and vision tasks.
+To emphasize the effectiveness of the proposed approach, an in-depth comparison of all these models has been conducted, achieving significant improvements in various evaluation metrics.
+IV-E Experiment metrics
 Used four metrics, Accuracy [46, 47, 48], F1 Score [49], IBA [50], and GMean [51], to assess the performance of the model holistically.
 The F1 Score is a crucial metric that combines accuracy and recall, providing a comprehensive measure of a model’s ability to classify positive and negative instances.
-This is especially useful for imbalanced datasets where positive and negative samples differ significantly.
-[52].
+This is especially useful for imbalanced datasets where positive and negative samples differ significantly [52].
 IBA (Imbalance Balanced Accuracy) takes a targeted approach to tackle the class imbalance challenge by harmonizing accuracy and balance.
@@
 GMean (Geometric Mean) is a metric that calculates the geometric mean of individual classification accuracies, which is used to evaluate the model’s ability to handle category imbalance and ensure that the model can perform well across all classes [54].
-In Table 5 showcases the performance of the model on the test set across various image resolutions, such as 40X, 100X, 200X, and 400X.
+Table 6 showcases the performance of the model on the test is set across various image resolutions, such as 40X, 100X, 200X, and 400X.
 Regardless of the solution, the model consistently outperforms nine other popular network models regarding accuracy, F1 score, IBA, and GMean.
@@
 This gives us a deep insight into the model’s excellent performance and great potential in performing medical image classification tasks.
-In Table 6, presents a comparison of the model with eight other network models regarding the number of model parameters and convergence speed.
+Table 7 compares the model with eight other network models regarding the number of model parameters and convergence speed.
 Notably, the DenseNet-121 model possesses the fewest parameters.
@@
 These results proved the efficiency and effectiveness of the model.
-5 Performance evaluation and visualization
+V Performance evaluation and visualization
 To thoroughly oversee the model’s performance and glean insights into the training process, focus on several key metrics, including training loss, validation loss, training accuracy, and validation accuracy.
@@
 Fig. 7: Training and validation loss with error bars
-5.1 Loss analysis
+Refer to caption 40X
+Refer to caption 100X
+Refer to caption 200X
+Refer to caption 400X
+Fig. 8: Training and validation accuracy with error bars
+V-A Loss analysis
 This study utilized a dual-activated lightweight attention ResNet50 model for breast cancer classification.
@@
 The x-axis denotes the quantity of training iterations, while the y-axis signifies the values of the loss function.
-Used line plots to show the trend of training loss and validation loss and added error bars to each data point to show the standard deviation of the loss values.
+Line plots were used to show the trend of training loss and validation loss, and error bars for each data point were used to indicate the standard deviation of the loss values.
 As training progresses, a gradual decrease in training loss and validation loss is observed, which signifies that the model is effectively learning and adapting to the data.
@@
 It offers a stable trend, indicating that the model steers from overfitting on the validation set and demonstrates strong generalization to new data.
-Refer to caption 40X
-Refer to caption 100X
-Refer to caption 200X
-Refer to caption 400X
-Fig. 7: Training and validation accuracy with error bars
-5.2 Accuracy analysis
+V-B Accuracy analysis
 In Fig.
-7, the evolution of training accuracy and validation accuracy at different image resolutions (40X, 100X, 200X, and 400X) is depicted [58].
+8, the evolution of training accuracy and validation accuracy at different image resolutions (40X, 100X, 200X, and 400X) is depicted [58].
 The horizontal axis represents the number of training iterations, while the vertical axis represents the accuracy.
@@
 These findings are pivotal in comprehending the model’s performance and can facilitate further optimization for enhanced breast cancer classification results.
-6 Summary and conclusions
+VI Summary and conclusions
 This paper focuses on applying a lightweight attention mechanism to a ResNet50 model for classifying medical histopathology tissue images, mainly when dealing with limited-scale datasets.
@@
 These findings provide a solid basis for enhancing classification accuracy.
-In future research, investigating more breast multiclassification problems is planned.
-[…]
+In future research, investigating more investigations of breast multiclassification problems are planned.

### v8

@@
 Abstract
-Automatic breast cancer classification technology in histopathology images plays a crucial role in accurate diagnosis and treatment planning.
-Recently, histopathology image classification approaches based on ResNet architecture have been widely proposed as they vastly improve the classification accuracy using skip connection to prevent vanishing gradient issues by combining low-level and high-level feature information.
-However, traditional ResNet architecture suffers from challenges such as data imbalance and interpretability, which still require cross-domain knowledge and collaboration among medical experts.
-This study proposes a new breast cancer classification method using a Dual-Activated Lightweight Attention ResNet50 (DALAResNet50) approach to address data imbalance and interpretability challenges effectively.
-The proposed approach combines a pre-trained deep ResNet50 method and a lightweight attention mechanism to accomplish the classification by embedding an attention module in layer 4 of ResNet50 and adding two fully connected layers based on LeakyReLU and ReLU activation functions to enhance the ability of feature learning.
-The proposed DALAResNet50 approach has been evaluated on breast cancer histopathology images derived from BreakHis Datase at magnification factors of 40X, 100X, 200X, and 400X, achieving accuracies of 98.5%, 98.7%, 97.9%, and 94.3%, respectively.
-The DALAResNet50 approach was comprehensively assessed with popular deep learning approaches such as SEResNet50, DensNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50 models.
-The reported results of DALAResNet50 outperform the compared approaches regarding precision, accuracy, recall, F1 score, and GMean, demonstrating significant robustness and broad applicability when dealing with different magnifications and unbalanced breast cancer datasets.
+Automatic breast cancer classification in histopathology images is crucial for precise diagnosis and treatment planning.
+Recently, classification approaches based on the ResNet architecture have gained popularity for significantly improving accuracy by using skip connections to mitigate vanishing gradient problems, thereby integrating low-level and high-level feature information.
+Nevertheless, the conventional ResNet architecture faces challenges such as data imbalance and limited interpretability, necessitating cross-domain knowledge and collaboration among medical experts.
+This study effectively addresses these challenges by introducing a novel method for breast cancer classification, the Dual-Activated Lightweight Attention ResNet50 (DALAResNet50) model.
+It integrates a pre-trained ResNet50 model with a lightweight attention mechanism, embedding an attention module in the fourth layer of ResNet50 and incorporating two fully connected layers with LeakyReLU and ReLU activation functions to enhance feature learning capabilities.
+The DALAResNet50 method was tested on breast cancer histopathology images from the BreakHis Database across magnification factors of 40X, 100X, 200X, and 400X, achieving accuracies of 98.5%, 98.7%, 97.9%, and 94.3%, respectively.
+It was also compared with established deep learning models such as SEResNet50, DenseNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50.
+The reported results of DALAResNet50 have been shown to outperform the compared approaches regarding accuracy, F1 score, IBA, and GMean, demonstrating significant robustness and broad applicability when dealing with different magnifications and imbalanced breast cancer datasets.
 Index Terms:
@@
 I Introduction
-Breast cancer, particularly among women, has emerged as one of the leading causes of fatal diseases, with its incidence on a continuous rise, eliciting worldwide health concerns [1, 2].
-Early and accurate diagnosis and treatment are crucial in reducing the mortality rate of breast cancer.
-However, medical practices grapple with several obstacles, including inadequate healthcare resources and human error in interpretation, which hampers accurate and prompt detection and diagnosis of breast cancer.
-In this context, intelligent assisted diagnosis technology has become an innovative solution in breast cancer.
-Recent advancements in deep learning have had a transformative impact on breast cancer diagnosis.
-Deep learning theory has not only surpassed the performance limitations of traditional classification and recognition algorithms but has also integrated the entire feature extraction, feature selection, and classification process into a single step.
-Deep learning has simplified the training process by automatically extracting and classifying features through neural networks, eliminating redundancy and improving efficiency.
-These developments have revolutionized breast cancer diagnosis by improving accuracy and streamlining workflows [3].
-This profound learning advancement has revolutionized breast cancer diagnosis [4, 5, 6, 7].
-One key innovation in this area is developing and applying Convolutional Neural Networks (CNNs), a specific class of deep learning models designed to process data, such as medical images, with a grid-like structure.
-The concept of CNN was first introduced back in 1998 [8].
-The concept describes a method for efficiently classifying image data.
-CNNs gradually extract abstract features of an image through multi-level convolutional and pooling operations, allowing the network to remove necessary features from different categories of pictures so that the other categories can be classified accurately based on those features.
-CNNs have three main layers: convolutional, pooling, and fully connected.
-These layers work together to make CNNs a powerful tool for classifying medical images.
-The convolutional layer captures the image’s local features, the pooling layer reduces data dimensions while preserving significant information, and the fully connected layer maps these extracted features to various output classes, As shown in Fig.
+Breast cancer has become a leading cause of mortality, particularly among women, with its incidence rising continuously, raising global health concerns [1, 2].
+Early, accurate diagnosis and treatment are pivotal in decreasing the mortality rate associated with breast cancer.
+Yet, medical practices face challenges such as insufficient healthcare resources and human error in interpretation, which impede the timely and precise detection and diagnosis of breast cancer.
+Against this backdrop, intelligent diagnostic assistance technologies have emerged as innovative solutions in the battle against breast cancer.
+Recent advances in deep learning have revolutionized breast cancer diagnosis, surpassing the limitations of traditional classification and recognition algorithms by integrating feature extraction, selection, and classification into a unified step.
+Deep learning automates feature extraction and classification through neural networks, reducing redundancy and enhancing efficiency, thus significantly improving diagnostic accuracy and workflow [3, 4, 5, 6, 7].
+A pivotal innovation in this domain is developing and applying Convolutional Neural Networks (CNNs), specialized deep learning models adept at processing data, such as medical images, with a grid-like topology.
+In 1998, LeCun et al.
+[8] introduced Convolutional Neural Networks (CNNs), which progressively extract abstract features from images through multiple levels of convolutional and pooling operations.
+This process enables CNNs to achieve classification based on these extracted features.
+CNNs comprise convolutional, pooling, and fully connected layers, collaboratively facilitating the robust classification of medical images.
+The convolutional layer identifies local image features; the pooling layer reduces data dimensions while retaining critical information; and the fully connected layer maps these features to output classes, as illustrated in Fig.
 1.
-However, despite the considerable achievements of deep learning in medical image classification, conventional CNN architectures struggle with unbalanced medical image data.
-To tackle this issue, In this paper, a lightweight attention mechanism based on ResNet50 [9] is introduced to address the challenge of unbalanced medical image data in conventional CNN architectures.
-This novel approach effectively handles unstable category distributions and accurately detects malignant breast cancer in medical images.
-The critical contributions of the work can be summarized as follows:
-(a) Proposing a lightweight attention mechanism network that adaptively adjusts channel feature extraction to improve feature representation.
-(b) Integrating the attention mechanism module and ResNet50 to form a feature-rich model, the model achieves more efficient feature extraction based on fusing the two techniques.
-Specifically, in the fully connected network component of the attention mechanism module, LeakyReLU [10] and ReLU [11] activation functions were utilized.
-LeakyReLU was selected for its ability to mitigate the gradient vanishing problem.
-In contrast, ReLU was employed to enhance the representation of image features.
-This choice was instrumental in better capturing multi-scale spatial patterns.
-(c) In this paper, test images with different magnifications and verify that the network is invariant to image scales and performs better than other state-of-the-art methods.
-(d) The model was thoroughly evaluated by analyzing loss.
-The results show stable performance in the model’s later training phases, minimizing overfitting and exhibiting strong generalization for unfamiliar data.
-This paper is organized as follows:
-Section 2 provides an overview of the current state of the art in applying deep learning to breast cancer diagnostic techniques.
-Section 3 details the proposed breast cancer classification method based on dual activation lightweight attention ResNet50.
-Section 4 presents the dataset and implementation details of the experimental task and compares the classification results with other state-of-the-art networks.
-Section 5 comprehensively evaluates the model’s performance through loss and accuracy analysis.
-Finally, Section 6 provides a summary and conclusions.
+Despite deep learning’s significant achievements in medical image classification, conventional CNN architectures often struggle with imbalanced medical image data.
+This paper proposed a Dual-Activated Lightweight Attention ResNet50 (DALAResNet50) model designed to overcome the challenge of imbalanced data in CNN architectures.
+This innovative approach efficiently manages uneven category distributions and enhances the detection of malignant breast cancer in medical images.
+The primary contributions of this work are summarized as follows:
+(a) This study introduces a novel lightweight attention mechanism network that adaptively adjusts channel feature extraction to enhance feature representation.
+This approach optimizes the network’s ability to discern relevant features for breast cancer classification.
+(b) Integrating the attention mechanism with ResNet50 [9] creates a feature-rich model, achieving more efficient feature extraction by combining the strengths of both techniques.
+Specifically, within the attention mechanism’s fully connected network component, LeakyReLU [10] and ReLU [11] activation functions are employed.
+LeakyReLU is chosen for its capability to alleviate the issue of vanishing gradients; ReLU is used to improve the representation of image features, facilitating the capture of multi-scale spatial patterns more effectively.
+(c) The paper tests images at various magnifications, demonstrating the network’s scale invariance and superior performance over other contemporary methods.
+(d) The model undergoes comprehensive evaluation, focusing on analyzing loss.
+The findings indicate stable performance during the later stages of training, effectively minimizing overfitting and demonstrating strong generalization capabilities for new data.
+Organization of the paper:
+Section II reviews the latest advancements in applying deep learning for breast cancer diagnosis.
+Section III proposed the DALAResNet50 model approach for classifying breast cancer histopathology images.
+Section IV outlines the dataset and implementation specifics of the experimental setup and compares the classification outcomes with those of other leading networks.
+Section V conducts an in-depth performance evaluation of the model through loss and accuracy analysis.
+Section VI provides a summary and conclusions.
 II Related works
-Currently, two main methods are primarily used for cancer tissue histopathology image classification: traditional manual feature extraction and deep learning-based methods.
-Traditional methods rely on manually extracting image features and then using classifiers such as SVM or Random Forest for recognition; however, this method has shortcomings, such as a time-consuming feature extraction process, the requirement of domain knowledge, and the challenge of securing high-quality features [12].
-In contrast, methods based on deep learning extract features through models like convolutional neural networks, which is more efficient [13].
-Deep learning-based cancer tissue histopathology image classification primarily encompasses three methods: CNN-based breast cancer classification, transfer learning-based breast cancer classification, and attention mechanism-based breast cancer classification.
-The following section delves into these techniques in detail.
+Cancer tissue histopathology image classification predominantly utilizes traditional manual feature extraction and deep learning-based methods.
+Traditional approaches depend on manually extracting features from images and employing classifiers such as SVM (Support Vector Machine) or Random Forest for recognition.
+However, these methods encounter limitations, such as the need for intensive feature extraction, a requirement for substantial domain knowledge, and challenges in acquiring high-quality features [12].
+In contrast, deep learning-based methods, mainly through models like Convolutional Neural Networks (CNNs), offer a more efficient feature extraction process [13].
+Specifically, deep learning approaches to cancer tissue histopathology image classification include CNN-based, transfer learning-based, and attention mechanism-based breast cancer classification.
+Each of these methods leverages the inherent strengths of deep learning to improve classification performance, which will be explored in detail in the subsequent section.
 II-A CNN
-CNN [8], one of the most prevalent deep learning techniques, is apt for tasks involving image processing [14].
-Through convolutional, pooling, and fully connected layers, CNNs are highly effective in learning image features and classifying them [15].
-In breast cancer histopathology image classification, CNN models with various structures, such as AlexNet, Inception, VGGNet, DenseNet, ResNet, etc., can be trained on large datasets to learn features and classify breast cancer.
-AlexNet, a deep convolutional neural network introduced by Alex et al.
-in 2012 [16], marked a pivotal breakthrough in the ImageNet image classification competition, setting a landmark in deep learning and CNNs.
-Omonigho et al.
-[17] employed a modified AlexNet with augmentation techniques on the MIAS mammography database for breast cancer detection, achieving an impressive 95.70% accuracy.
-This surpassed traditional methods, though parameter optimization and validation on unbalanced datasets require further exploration.
-The Inception model is a neural network architecture introduced by Google that aims to solve the computation problem and several parameters in neural networks and is suitable for image classification and related tasks.
-The hybrid model based on CNN-Inception-V4 was used by Nazir et al.
-[18].
-It has excellent accuracy and sensitivity, but the model’s specificity is a little inferior, and it does not consider the processing computational efficiency and resource consumption issues.
-VGGNet, a deep neural network architecture developed by researchers at Oxford University, performs well in image classification tasks.
-However, it is unsuitable for embedded devices or resource-constrained environments because of its high computational and memory requirements due to its considerable depth and number of parameters.
-Nedjar et al .[19] proposed a visualization method for explaining and understanding the decision-making process of VGG19 in classifying breast biopsy images, which helps to increase the trust and acceptance of medical professionals while providing a method for evaluating and Enhancing the efficiency of CNNs.
-DenseNet is a deep neural network architecture developed by Gao Huang et al.
-[20].
-It is known for its unique connectivity pattern.
-Each layer is densely connected to all previous layers, thus enabling the effectiveness of information flow and gradient propagation in deep networks.
+CNN, one of the most widespread deep learning techniques, excels in image processing tasks [14].
+They utilize convolutional, pooling, and fully connected layers to learn and classify image features effectively [15].
+In breast cancer histopathology image classification, CNN models like AlexNet, Inception, VGGNet, DenseNet, and ResNet have demonstrated substantial capabilities when trained on large datasets.
+AlexNet, introduced in 2012 by Alex et al.
+[16], represented a significant milestone in ImageNet competition history, showcasing the potent capabilities of CNNs.
+Modified versions of AlexNet, such as the one by Omonigho et al.
+[17] on the MIAS mammography database, achieved notable accuracy, indicating superior performance over traditional methods.
+However, challenges remain in parameter optimization and application to imbalanced datasets.
+The Inception model, developed by Google, addresses computational and parameter efficiency in neural networks, making it ideal for image classification tasks.
+Nazir et al.’s hybrid CNN-Inception-V4 model [18] demonstrated high accuracy and sensitivity, though it faced limitations in specificity and computational efficiency considerations.
+VGGNet, crafted by Oxford University researchers, excels in image classification but is less suited for resource-constrained environments due to its intensive computational and memory demands.
+Nedjar et al.
+[19] presented a visualization method for VGG19’s decision-making process in breast biopsy image classification, enhancing medical professionals’ trust and CNN efficiency.
+DenseNet [20], known for its efficient information and gradient flow through dense connectivity, offers advantages in overcoming challenges like gradient vanishing.
 Jiménez Gaona et al.
-[21] introduced a novel approach to breast cancer detection and classification using the DenseNet architecture, which has advantages over other pre-trained CNN methods in solving problems such as gradient vanishing and reducing the number of parameters.
-ResNet, a deep convolutional neural network architecture introduced by He et al.
-in 2015 [9], is particularly notable for its solution to the vanishing gradient issue in deep neural networks.
-This enables the development of intense neural networks, leading to substantial improvements in tasks like image categorization.
-Khikani et al.
-[22] proposed an enhanced capsule network using Res2Net blocks and additional convolutional layers to extract multi-scale features, and the results showed that their model outperformed previous deep-learning methods for breast cancer classification.
-However, ResNet-50 is prone to problems in handling unbalanced datasets.
-Given the significant superiority of the ResNet50 model over other methods in terms of accuracy, this paper has decided to modify the ResNet50 model to address these specific challenges.
+[21] leveraged DenseNet for breast cancer detection, showing its superiority in reducing parameters and improving performance compared to other CNN architectures.
+ResNet, introduced by He et al.
+in 2015 [9], addresses deep neural networks’ vanishing gradient problem, allowing for the construction of deeper networks.
+An enhanced capsule network using Res2Net blocks, proposed by Khikani et al.
+[22], demonstrated improved performance in breast cancer classification, although ResNet-50 faces challenges with imbalanced datasets.
+This paper presents modifications to the ResNet50 model to tackle these issues, given its accuracy advantage over other methods.
 II-B Transfer learning
@@
 II-C Attention mechanism
-In medical imaging, the attention mechanism [32] in deep learning has proven to be a groundbreaking technique.
-It enables models to assign varying levels of importance to different regions within medical images, such as MRI scans, X-rays, CT scans, and histopathology images [33].
-This targeted focus allows the models to prioritize and analyze the most pertinent areas for diagnosing diseases, identifying anomalies, or tracking the progression of medical conditions.
-By concentrating on crucial features within these images, the attention mechanism dramatically enhances the accuracy and efficiency of automated medical image analysis, aiding in early detection and improved patient outcomes.
-In the context of breast cancer image classification, introducing the attention mechanism approach can prompt the model to learn critical and discriminative features in breast cancer images crucial for accurately distinguishing between malignant and benign tissues.
-This technique can improve the model’s performance while reducing the reliance on global image feature extraction and improving the accuracy of identifying critical breast cancer regions [34].
+In medical imaging, the attention mechanism [32] in deep learning has emerged as a revolutionary technique.
+It allows models to allocate varying levels of importance to different regions within medical images, such as MRI scans, X-rays, CT scans, and histopathology images [33].
+This selective focus enables the models to prioritize and scrutinize the most relevant areas for diagnosing diseases, detecting anomalies, or monitoring the progression of medical conditions.
+By zeroing in on essential features within these images, the attention mechanism significantly boosts the accuracy and efficiency of automated medical image analysis, contributing to early detection and enhanced patient outcomes.
+Within breast cancer image classification, incorporating the attention mechanism encourages the model to discern critical and distinctive features in breast cancer images that are pivotal for accurately differentiating between malignant and benign tissues.
+This strategy can elevate the model’s performance, diminish reliance on global image feature extraction, and enhance the precision in pinpointing vital breast cancer regions [34].
 Researchers have proposed various breast cancer segmentation and classification systems incorporating attention mechanisms.
-For instance, Vardhan et al.
-[35] developed a breast cancer segmentation system that uses convolutional neural networks with interpretable artificial intelligence and the attention mechanism to improve breast cancer detection and classification accuracy and speed.
+For example, Vardhan et al.
+[35] developed a breast cancer segmentation system utilizing Convolutional Neural Networks (CNNs) with interpretable artificial intelligence and the attention mechanism to enhance breast cancer detection and classification accuracy and speed.
 Xu et al.
-[36] designed a system for breast cancer segmentation that uses a modified DenseNet architecture with an attention mechanism to optimize the convolutional neural network on the BreakHis dataset for high-precision classification of breast cancer.
-The method was tested on the BreakHis dataset, and the method’s performance was evaluated using accuracy metrics for both binary and multi-classification.
+[36] designed a system for breast cancer segmentation employing a modified DenseNet architecture with an attention mechanism to refine the CNN’s performance on the BreakHis dataset for high-precision breast cancer classification.
+This method was evaluated on the BreakHis dataset using accuracy metrics for both binary and multi-classification.
 Wang et al.
-[37] proposed a breast cancer classification model that combines deep migration learning and visual attention mechanisms to improve the accuracy and efficiency of real-world clinical diagnosis.
-This deep-migration-based attention mechanisms approach uses migration learning to overcome the overfitting problem and combines the attention module to focus on characterizing breast lesions.
-Wang et al.’s work [37] achieves high-precision results that outperform other state-of-the-art breast image classification algorithms.
-This section summarizes recent scholarly research and exploration in breast cancer classification, focusing on the use of the attention mechanism.
-The insights gleaned indicate that applying the attention mechanism in breast cancer images and classification methods can simultaneously enhance the accuracy of crucial region identification and boost the interpretability of the model.
+[37] proposed a breast cancer classification model that combines deep transfer learning and visual attention mechanisms to increase the accuracy and efficiency of real-world clinical diagnoses.
+Based on deep transfer learning with attention mechanisms, this approach utilizes transfer learning to tackle the overfitting problem and integrates an attention module to concentrate on characterizing breast lesions.
+Wang et al.’s work achieves high-precision results that surpass other state-of-the-art breast image classification algorithms.
+This section summarizes recent scholarly research and exploration of breast cancer classification, emphasizing the use of the attention mechanism.
+The insights indicate that applying the attention mechanism in breast cancer images and classification methods can simultaneously enhance the accuracy of crucial region identification and increase the model’s interpretability.
 However, this also correspondingly increases the following difficulties:
 II-C1 Limitations in global feature extraction
-Although the attention mechanism is designed to enhance model performance by focusing on critical parts of the image, thereby playing a positive role in strengthening the model’s ability to process global information, it may still face challenges in some implementations or specific application scenarios, unable to fully capture the international contextual details of the entire image.
+Although the attention mechanism is designed to improve model performance by focusing on critical sections of the image, thereby playing a positive role in enhancing the model’s ability to process global information of the image, it may still encounter challenges in some implementations or specific application scenarios, failing to grasp the entire image’s global contextual details fully.
 II-C2 Insufficiency in feature priority precision
 Lacking a sophisticated method to discern the significance of diverse features, attention mechanisms might uniformly weigh all channel information.
-This approach risks overlooking vital features essential for precise classification, undermining the model’s ability to highlight and utilize discriminative patterns within the data effectively.
+This approach risks overlooking vital features essential for precise classification, compromising the model’s effectiveness in identifying and leveraging distinctive patterns in the data.
 II-C3 Inefficiency of Activation Functions
@@
 These pioneering approaches synergistically aim to significantly elevate model performance while mitigating computational resource requirements, providing a more efficient and adaptable solution for integrating attention mechanisms into breast cancer image classification models.
+Moreover, this paper presents a comparative analysis between the proposed DALAResNet50 model and the standard ResNet model, as illustrated in Table 2.
 III Methods
@@
 Uniquely, the ResNet50 model introduces the concept of residual learning.
-This innovation solves the problem of vanishing gradients in deep networks, thereby opening up the possibility of building deeper neural networks than ever before.
-This design breakthrough allowed for achieving exceptional performance.
-However, considering the efficient utilization of data and computing resources, the transfer learning strategy is adopted in this paper based on the pre-trained ResNet50 model for model building.
-Through transfer learning, full benefit can be obtained from the high-level feature representations learned by ResNet50 on large-scale datasets in medical image classification tasks.
+This innovation solves the problem of vanishing gradients in deep networks, thereby opening up the possibility of building deeper neural networks.
+This design breakthrough allowed for achieving promising performance.
+Therefore, the transfer learning strategy based on the pre-trained ResNet50 model is adopted in this paper, considering the efficient utilization of data and computing resources for model building.
+The benefit of transfer learning is the high-level feature representations learned by ResNet50 on large-scale datasets in medical image classification tasks.
 This allows for swiftly attaining superior classification outcomes, even with constrained data and computing capacities.
+As shown in Fig.
+2, the proposed approach includes a histopathology breast cancer database and components for preprocessing and data augmentation, transfer learning based on ResNet50, and classification based on an improvised dual-activated fully connected layer.
+The proposed approach involves the following stages for breast cancer classification using histopathology images:
 TABLE I: The ResNet50 Model
@@
 Pooling $1\times 11 × 1$ Average Pool, 1000-d FC.
-Note: The term ”1000-d FC” in the pooling layer refers to a fully-connected layer consisting of 1000 neurons.
+Note: The term ”1000-d FC” in the pooling layer refers to a fully connected layer consisting of 1000 neurons.
 TABLE II: Comparison between DALAResNet50 model and standard ResNet model
-Feature
-DALAResNet50 model
-Standard ResNet model
-Adaptive Average Pooling
-Adaptive average pooling addresses ”limitations in global feature extraction” by enhancing the model’s capacity to process global contextual information, significantly improving performance in tasks requiring comprehensive image understanding.
+Feature DALAResNet50 model Standard ResNet model
+Adaptive Average Pooling Adaptive average pooling addresses ”limitations in global feature extraction” by enhancing the model’s capacity to process global contextual information, significantly improving performance in tasks requiring comprehensive image understanding.
 Not available
-Channel Recalibration
-A two-layer fully connected network for channel recalibration optimizes ”feature priority precision,” enhancing the model’s ability to identify and focus on the most relevant features for accurate classification.
+Channel Recalibration A two-layer fully connected network for channel recalibration optimize ”feature priority precision,” enhancing the model’s ability to identify and focus on the most relevant features for accurate classification.
 Not available
-Activation Functions
-The dual activation functions of LeakyReLU and ReLU address the ”inefficiency of activation functions.” By incorporating LeakyReLU, the network avoids the dying ReLU problem, as it allows a small, non-zero gradient when the unit is not active, ensuring continuous learning dynamics and preventing neurons from becoming inactive across all network layers.
+Activation Functions The dual activation functions of LeakyReLU and ReLU address the ”inefficiency of activation functions.” By incorporating LeakyReLU, the network avoids the dying ReLU problem, as it allows a small, nonzero gradient when the unit is not active, ensuring continuous learning dynamics and preventing neurons from becoming inactive across all network layers.
 This combination maintains the model’s learning capability, potentially enhancing the training process.
 ReLU only
-Element-wise multiplication
-Element-wise multiplication, as denoted by ’Scale’ in Fig.
-3, enhances ”Feature Adaptability” by allowing dynamic, non-linear adjustment of features according to a specific input, improving the network’s flexibility and utilization of information.
+Element-wise multiplication Element-wise multiplication, as denoted by ’Scale’ in Fig.
+3, enhance ”Feature Adaptability” by allowing dynamic, nonlinear adjustment of features according to a specific input, improving the network’s flexibility and utilization of information.
 Not available
-Fig. 2: Flowchart for standard ResNet model
-Fig. 3: Flowchart for DALAResNet50 model
+Fig. 2: The framework of DALAResNet50 model
+Fig. 3: Comparison of conventional ResNet50 and DALAResNet50 model: (a) ResNet50 model (b) DALAResNet50 model
 Fig. 4: Visualization of DALAResNet50 model
-Fig. 5: The framework of DALAResNet50 model
 III-B Dual-activated lightweight attention ResNet50
 A new approach for automatic breast cancer classification is proposed in this research using a new version of the ResNet50 technique called Dual-activated lightweight attention ResNet50 (DALAResNet50).
-The DALAResNet50 approach incorporates ResNet50, Global pooling, channel recalibration, Dual Activation functions, and the newly designed fully connected layer into a unified approach to overcome the drawbacks of the traditional ResNet50 model.
-The pre-trained ResNet50 is first utilized as a baseline architecture.
-Subsequently, the attention module is integrated into the fourth residual block of ResNet50 to pay attention to the small-scale nuclei cells of breast cancer.
-The attention mechanism improves the ability of feature learning by extracting the silent semantic information to overcome the class imbalance issue.
-The proposed attention module is characterized by the amalgamation of adaptive global pooling and a two-layer fully connected network, employing LeakyReLU [10] and ReLU [11] as dual activation functions.
-The traditional ResNet50 uses conventional pooling layers to reduce the spatial dimensions of the feature maps while retaining important information.
-However, the conventional pooling layers operate over local regions of the input feature maps, reducing their spatial dimensions but discarding spatial information, making the ResNet50 sensitive to spatial translations of the input feature map and increasing the number of parameters in the network.
-Therefore, this research proposed average-based global pooling that calculates the average value of each feature map across all spatial locations, effectively summarizing each feature’s presence throughout the entire feature map.
-The main advantage of global pooling is that it provides a fixed-size representation of the whole feature map, making the ResNet50 less sensitive to spatial translations of the input and reducing the total number of parameters in the network.
-Additionally, the DALAResNet50 adopts the LeakyReLU activation function [10] for the initial fully connected layer, leveraging its competence in allowing for a slight gradient flow under zero.
-This is done to mitigate the issue of vanishing gradients, thereby enhancing the training efficacy of deep neural networks.
-The ReLU activation function [11] is further chosen for the second fully connected network to improve the model performance and image feature representation.
-Combining these two activation functions enhances the model’s expressive power and capacity to learn intricate image features, making the model more suitable for breast cancer image classification tasks.
-Finally, the newly added fully connected layer is positioned at the end of the DALAResNet50 approach to address the complexity of binary classification problems.
+The DALAResNet50 model is methodically structured into three core stages: data preprocessing and augmentation, transfer learning based on ResNet50 architecture, and DALAResNet50 model optimization, as detailed below:
+III-B1 Data preprocessing and augmentation
+Initially, all breast cancer histopathology images are resized to 224x224 pixels, aligning with the input specifications for convolutional neural networks.
+A normalization process follows, adjusting pixel values to a range between [-1, 1] to expedite model training and augment the model’s generalization capability with novel data.
+Data augmentation strategies like random rotation and horizontal flipping bolster the model’s robustness.
+These strategies mimic potential real-world image alterations, improving the model’s resilience to subtle image variations.
+III-B2 Transfer learning based on ResNet50
+At this stage, capitalizes on a pre-trained ResNet50 framework, incorporating a custom-designed attention module within its fourth residual block.
+This module aims to pinpoint small-scale cell nuclei characteristic of breast cancer and glean pertinent semantic information.
+This strategic enhancement significantly boosts the model’s ability to learn and discern features.
+III-B3 DALAResNet50 model optimization
+During the optimization phase of the DALAResNet50 model, it is adopted that focuses sequentially on pooling layers, activation layers, and nonlinear feature adjustments to enhance the model’s performance and adaptability.
+The related steps are as follows:
+  1. 1.
+    Initially, to address the issue of spatial variations in input feature maps, traditional pooling layers are replaced with adaptive average pooling.
+This adjustment optimizes the network’s parameter structure, enabling the model to process global contextual information more effectively and accurately.
+The introduction of adaptive average pooling not only simplifies the network’s complexity but also enhances the model’s ability to capture important information from different parts of the image, laying a solid foundation for subsequent feature processing.
+  2. 2.
+    Next, the model employs a dual activation function strategy in the activation layer, starting with the LeakyReLU activation function [10] in the initial phase.
+This function is chosen for its unique capability to maintain a slight yet nonzero gradient when the unit is inactive, effectively combating the problem of vanishing gradients and significantly boosting the training efficiency of deep neural networks.
+Following this, the ReLU activation function [11] is applied in the subsequent fully connected layer to optimize model performance further and enhance the representation of image features.
+The careful combination of LeakyReLU and ReLU activation functions dramatically improves the model’s ability to process and learn complex image characteristics, providing strong support for accurately classifying breast cancer images.
+  3. 3.
+    Finally, the introduction of element-wise multiplication operations further enhances the model’s Feature Adaptability, allowing for dynamic, nonlinear adjustments of features based on specific inputs.
+This step significantly increases the network’s flexibility and information utilization, enabling the model to adapt its processing flow according to the particular circumstances of the input features.
+This approach notably improves the model’s feature processing capability, leading to higher accuracy and efficiency in classifying breast cancer images.
 Fig.
-2 and Fig.
-3 illustrate the flowcharts of the standard ResNet approach and the DALAResNet50 approach, respectively.
+3 illustrates the flowcharts of the standard ResNet50 approach and the DALAResNet50 approach, respectively.
 These flowcharts highlight the newly added features in the DALAResNet50.
 A comprehensive comparative analysis of these differences is detailed in Table 2.
-The visualization of the DALAResNet50 approach and the entire technical framework are depicted in Fig.
-4 and Fig.
-5.
+The visualization of the DALAResNet50 approach is depicted in Fig.
+4.
 IV Experiment
 IV-A BreakHis Dataset
-The BreakHis dataset, utilized in this study focusing on breast cancer histopathological images, was generously provided by scholar Spanha [38].
+The BreakHis dataset, utilized in this study focusing on breast cancer histopathology images, was generously provided by scholar Spanha [38].
 This dataset comprises histopathology samples of breast cancer tumors obtained from 82 patients.
@@
 5 presents these pathological images as a practical example.
-Refer to caption 40x
-Refer to caption 100x
-Refer to caption 200x
-Refer to caption 400x
+40x
+100x
+200x
+400x
 Fig. 5: Images of breast cancer pathology at four magniations
@@
 40X 100X 200X 400X
-Benign
-625 644 623 588
-Malignant
-1370 1437 1390 1232
-Total No of Images
-1995 2081 2013 1820
+Benign 625 644 623 588
+Malignant 1370 1437 1390 1232
+Total No of Images 1995 2081 2013 1820
 TABLE IV: Dataset split for experimental task
@@
 Considering the relatively small size of the BreakHis medical imaging dataset (containing 7909 images) and the uneven distribution of the data, a lower learning rate of 0.0001 was chosen for this study so that the model could adjust the weights more accurately.
-This is important for dealing with small and unbalanced datasets to ensure that all categories are given proper attention.
+This is important for dealing with small and imbalanced datasets to ensure that all categories are given proper attention.
 In addition, setting the training batch size and the number of iterations is significant.
@@
 Previous studies generally set the dropout rate between 0.25 and 0.5.
-In processing the unbalanced BreakHis dataset, a lower dropout rate helps ensure that the categories with smaller numbers are sufficiently learned, thus improving the model’s recognition accuracy and generalization ability.
+In processing the imbalanced BreakHis dataset, a lower dropout rate helps ensure that the categories with smaller numbers are sufficiently learned, thus improving the model’s recognition accuracy and generalization ability.
 Therefore, this study chose a lower dropout rate of 0.25.
-The method proposed in this study, and the benchmark model illustrated in the Table apply the above settings to ensure the experiments’ fairness and the comparability of the results.
+The method proposed in this study and the benchmark model illustrated in Table 6 apply the above settings to ensure the experiments’ fairness and the comparability of the results.
 These carefully selected parameters provide a solid foundation for the experimental design and ensure the reliability and efficiency of the model throughout the study; the parameter settings are shown in Table 5.
@@
 Accuracy F1 IBA GMean Accuracy F1 IBA GMean
-SEResNet50 [39]
-0.883 0.779/0.883 0.860 0.860 0.850 0.720/0.849 0.813 0.812
-DensNet121 [40]
-0.801 0.639/0.800 0.763 0.762 0.811 0.654/0.809 0.755 0.753
-VGG16 [41]
-0.902 0.813/0.902 0.877 0.877 0.850 0.724/0.851 0.830 0.829
-VGG16Inception [42]
-0.977 0.955/0.977 0.965 0.965 0.971 0.943/0.971 0.964 0.964
-ViT [43]
-0.835 0.693/0.833 0.769 0.767 0.834 0.694 /0.833 0.788 0.787
-Swin-Transformer [44]
-0.843 0.711/0.844 0.819 0.818 0.877 0.767/0.876 0.836 0.835
-Dinov2_Vitb14 [45]
-0.873 0.762/0.873 0.843 0.843 0.869 0.753/0.868 0.835 0.834
-RseNet50 [9]
-0.927 0.860/0.927 0.914 0.914 0.944 0.891/0.944 0.922 0.922
+SEResNet50 [39] 0.883 0.779/0.883 0.860 0.860 0.850 0.720/0.849 0.813 0.812
+DensNet121 [40] 0.801 0.639/0.800 0.763 0.762 0.811 0.654/0.809 0.755 0.753
+VGG16 [41] 0.902 0.813/0.902 0.877 0.877 0.850 0.724/0.851 0.830 0.829
+VGG16Inception [42] 0.977 0.955/0.977 0.965 0.965 0.971 0.943/0.971 0.964 0.964
+ViT [43] 0.835 0.693/0.833 0.769 0.767 0.834 0.694 /0.833 0.788 0.787
+Swin-Transformer [44] 0.843 0.711/0.844 0.819 0.818 0.877 0.767/0.876 0.836 0.835
+Dinov2_Vitb14 [45] 0.873 0.762/0.873 0.843 0.843 0.869 0.753/0.868 0.835 0.834
+RseNet50 [9] 0.927 0.860/0.927 0.914 0.914 0.944 0.891/0.944 0.922 0.922
 DALAResNet50 Model 0.985 0.970/0.985 0.980 0.980 0.987 0.974/0.987 0.985 0.985
@@
 Accuracy F1 IBA GMean Accuracy F1 IBA GMean
-SEResNet50 [39]
-0.906 0.820/0.906 0.876 0.875 0.876 0.770/0.878 0.864 0.863
-DensNet121 [40]
-0.812 0.659/0.812 0.775 0.774 0.779 0.603/0.777 0.737 0.736
-VGG16 [41]
-0.851 0.722/0.850 0.812 0.811 0.848 0.718/0.848 0.817 0.817
-VGG16Inception [42]
-0.954 0.910/0.954 0.938 0.938 0.889 0.788/0.888 0.857 0.856
-ViT [43]
-0.841 0.705/0.840 0.801 0.800 0.832 0.691/0.831 0.803 0.803
-Swin-Transformer [44]
-0.876 0.766/0.875 0.838 0.837 0.846 0.724/0.851 0.791 0.789
-Dinov2_Vitb14 [45]
-0.853 0.726/0.852 0.802 0.800 0.885 0.782/0.885 0.840 0.839
-RseNet50 [9]
-0.960 0.922/0.960 0.947 0.947 0.925 0.856/0.925 0.918 0.918
+SEResNet50 [39] 0.906 0.820/0.906 0.876 0.875 0.876 0.770/0.878 0.864 0.863
+DensNet121 [40] 0.812 0.659/0.812 0.775 0.774 0.779 0.603/0.777 0.737 0.736
+VGG16 [41] 0.851 0.722/0.850 0.812 0.811 0.848 0.718/0.848 0.817 0.817
+VGG16Inception [42] 0.954 0.910/0.954 0.938 0.938 0.889 0.788/0.888 0.857 0.856
+ViT [43] 0.841 0.705/0.840 0.801 0.800 0.832 0.691/0.831 0.803 0.803
+Swin-Transformer [44] 0.876 0.766/0.875 0.838 0.837 0.846 0.724/0.851 0.791 0.789
+Dinov2_Vitb14 [45] 0.853 0.726/0.852 0.802 0.800 0.885 0.782/0.885 0.840 0.839
+RseNet50 [9] 0.960 0.922/0.960 0.947 0.947 0.925 0.856/0.925 0.918 0.918
 DALAResNet50 Model 0.979 0.958/0.979 0.98 0.98 0.943 0.891/0.944 0.945 0.945
@@
 40X 100X 200X 400X
-SEResNet50 [39]
-2696.22 2814.30 2711.89 2356.18 27073090 2644.65
-DenseNet121 [40]
-2352.46 2643.79 2423.36 2078.34 9028970 2347.49
-VGG16 [41]
-3212.95 3182.54 3011.53 2640.19 134268738 3011.80
-VGG16Inception [42]
-3074.99 3253.50 3303.12 2682.61 38139938 3078.56
-ViT [43]
-2932.11 3022.54 3013.86 2691.16 51375106 2914.92
-Swin-Transformer [44]
-3020.11 3132.42 2940.98 2592.96 86745274 2921.62
-Dinov2_Vitb14 [45]
-2895.02 2303.25 2245.98 2017.15 86582018 2365.35
-ResNet50 [9]
-2920.53 2933.26 2307.94 2097.33 21797672 2564.77
+SEResNet50 [39] 2696.22 2814.30 2711.89 2356.18 27073090 2644.65
+DenseNet121 [40] 2352.46 2643.79 2423.36 2078.34 9028970 2347.49
+VGG16 [41] 3212.95 3182.54 3011.53 2640.19 134268738 3011.80
+VGG16Inception [42] 3074.99 3253.50 3303.12 2682.61 38139938 3078.56
+ViT [43] 2932.11 3022.54 3013.86 2691.16 51375106 2914.92
+Swin-Transformer [44] 3020.11 3132.42 2940.98 2592.96 86745274 2921.62
+Dinov2_Vitb14 [45] 2895.02 2303.25 2245.98 2017.15 86582018 2365.35
+ResNet50 [9] 2920.53 2933.26 2307.94 2097.33 21797672 2564.77
 DALAResNet50 Model 2133.02 2319.50 1990.58 1896.82 25611842 2084.98
@@
 Susan.
-The article delves into a deep learning model called ”VGGIN-Net,” which is mainly applied to analyze unbalanced breast cancer datasets.
+The article delves into a deep learning model called ”VGGIN-Net,” which is mainly applied to analyze imbalanced breast cancer datasets.
 The study introduces a new deep learning framework that incorporates transfer learning to address the problem of category imbalance in biomedical datasets, applied explicitly to the BreakHis breast cancer dataset.
-To reduce training time and computational cost, the model uses a pre-trained network to migrate learned weights from one domain to another.”
+The model uses a pre-trained network to migrate learned weights from one domain to another to reduce training time and computational cost.
 ViT (Vision Transformer) [43]: Proposed by Dosovitskiy et al.
@@
 To emphasize the effectiveness of the proposed approach, an in-depth comparison of all these models has been conducted, achieving significant improvements in various evaluation metrics.
+40X
+100X
+200X
+400X
+Fig. 6: Training and validation loss with error bars
 IV-E Experiment metrics
-Used four metrics, Accuracy [46, 47, 48], F1 Score [49], IBA [50], and GMean [51], to assess the performance of the model holistically.
-The F1 Score is a crucial metric that combines accuracy and recall, providing a comprehensive measure of a model’s ability to classify positive and negative instances.
+Used four metrics, Accuracy [46, 47, 48], F1 Score [49], Imbalance Balanced Accuracy (IBA) [50], and Geometric Mean (GMean) [51], to assess the performance of the model holistically.
+Below is an introduction to these metrics:
+  1. 1.
+    Accuracy is one of the most intuitive and commonly used metrics for evaluating the overall performance of a model.
+It calculates the proportion of all samples (whether positive or negative) that were predicted correctly relative to the total number of samples [46, 47, 48].
+  2. 2.
+    The F1 Score is a crucial metric that combines precision and recall, providing a comprehensive measure of a model’s ability to classify positive and negative instances.
 This is especially useful for imbalanced datasets where positive and negative samples differ significantly [52].
-IBA (Imbalance Balanced Accuracy) takes a targeted approach to tackle the class imbalance challenge by harmonizing accuracy and balance.
-This approach offers a more holistic assessment of the model’s performance, encompassing sample diversity and fairness [53].
-GMean (Geometric Mean) is a metric that calculates the geometric mean of individual classification accuracies, which is used to evaluate the model’s ability to handle category imbalance and ensure that the model can perform well across all classes [54].
-Table 6 showcases the performance of the model on the test is set across various image resolutions, such as 40X, 100X, 200X, and 400X.
-Regardless of the solution, the model consistently outperforms nine other popular network models regarding accuracy, F1 score, IBA, and GMean.
-These results demonstrate that the network model has a significant advantage regardless of which image resolution is used.
-This gives us a deep insight into the model’s excellent performance and great potential in performing medical image classification tasks.
-Table 7 compares the model with eight other network models regarding the number of model parameters and convergence speed.
-Notably, the DenseNet-121 model possesses the fewest parameters.
-However, the model has the shortest convergence time and the fastest average convergence speed across all image resolutions, including 40X, 100X, 200X, and 400X.
-This further reinforces the superior performance and potential of the model for medical image classification tasks.
-These results proved the efficiency and effectiveness of the model.
+  3. 3.
+    IBA is a metric for evaluating the performance of classification models on imbalanced datasets, adjusting standard accuracy to more fairly account for the model’s ability to identify both majority and minority classes, ensuring the performance assessment reflects classification effectiveness across all categories.
+[53].
+  4. 4.
+    GMean is a metric that calculates the geometric mean of individual class accuracies, indicating a model’s capability to address category imbalance.
+It effectively equalizes the performance assessment for both dominant and infrequent categories, emphasizing the model’s comprehensive efficiency in handling variances in class distribution.
+This metric is in scenarios where class imbalance could significantly skew the performance metrics, ensuring that the model’s ability to classify less represented classes correctly is adequately considered in its overall [54].
+Table 6 showcases the model’s performance on the test set at various image resolutions (40X, 100X, 200X, and 400X).
+Across these resolutions, the model consistently outperforms nine other popular networks regarding accuracy, F1 score, IBA, and GMean.
+This consistency underlines the network’s significant advantage and underscores its potential for medical image classification tasks.
+Table 7 compares the model against eight others regarding parameter count and convergence speed.
+While DenseNet-121 has the fewest parameters, the proposed model achieves the shortest convergence times, exhibiting the fastest average speed across all resolutions (40X, 100X, 200X, and 400X).
+This further reinforces the proposed model’s efficiency and strong potential for medical image classification, highlighting superior performance and effectiveness.
 V Performance evaluation and visualization
-To thoroughly oversee the model’s performance and glean insights into the training process, focus on several key metrics, including training loss, validation loss, training accuracy, and validation accuracy.
-Visualizing these metrics can demonstrate the model’s training performance.
-The results of analyzing these key metrics are presented in detail below.
-Refer to caption 40X
-Refer to caption 100X
-Refer to caption 200X
-Refer to caption 400X
-Fig. 7: Training and validation loss with error bars
-Refer to caption 40X
-Refer to caption 100X
-Refer to caption 200X
-Refer to caption 400X
+To effectively monitor the model’s performance and gain insights into the training process, it is essential to concentrate on various crucial metrics, such as training loss, validation loss, training accuracy, and validation accuracy.
+Visualizing these metrics offers a comprehensive view of the model’s training behavior.
+An in-depth analysis of these key metrics and their implications is outlined below:
+40X
+100X
+200X
+400X
 Fig. 7: Training and validation accuracy with error bars
 V-A Loss analysis
-This study utilized a dual-activated lightweight attention ResNet50 model for breast cancer classification.
-The critical metrics for monitoring model performance and understanding the training process are training loss and validation loss [55, 56, 57].
-These loss values play an important role when evaluating model performance, and their visualization gives us a clearer picture of the training progress.
+This study deployed a DALAResNet50 model for the classification of breast cancer.
+Critical metrics for assessing model performance and understanding the training dynamics include training loss and validation loss [55, 56, 57].
+These metrics are crucial for evaluating the model’s effectiveness, and their graphical representation offers a detailed insight into the training progression.
 Fig.
-7 shows the variation of training loss and validation loss with the number of training iterations for different image resolutions (including 40X, 100X, 200X, and 400X).
-The x-axis denotes the quantity of training iterations, while the y-axis signifies the values of the loss function.
-Line plots were used to show the trend of training loss and validation loss, and error bars for each data point were used to indicate the standard deviation of the loss values.
-As training progresses, a gradual decrease in training loss and validation loss is observed, which signifies that the model is effectively learning and adapting to the data.
-This is a positive trend and shows that the model is making progress on the task.
-In the later stages of training, the validation loss changes less.
-It offers a stable trend, indicating that the model steers from overfitting on the validation set and demonstrates strong generalization to new data.
+6 illustrates the fluctuations in training and validation losses across varying numbers of training iterations at different image magnifications (40X, 100X, 200X, and 400X).
+The x-axis represents the number of training iterations, while the y-axis reflects the loss values.
+Line plots demonstrate the trends in training and validation losses, with error bars on each point indicating the standard deviation of these values.
+Training and validation losses consistently decline, indicating the model’s successful learning and adaptation to the dataset.
+This downward trend is encouraging, as it signals ongoing improvement in task performance.
+In the later training phases, the validation loss stabilizes, suggesting that the model avoids overfitting on the validation dataset and exhibits robust generalization capabilities to unseen data.
 V-B Accuracy analysis
-In Fig.
-8, the evolution of training accuracy and validation accuracy at different image resolutions (40X, 100X, 200X, and 400X) is depicted [58].
-The horizontal axis represents the number of training iterations, while the vertical axis represents the accuracy.
-A line graph is employed to portray the accuracy trend, and error bars are incorporated to convey the standard deviation of the accuracy values.
+In Fig.7, the progression of training and validation accuracy across various image resolutions (40X, 100X, 200X, and 400X) is illustrated [58].
+The horizontal axis indicates the number of training iterations, whereas the vertical axis displays accuracy levels.
+A line graph is utilized to depict the trend in accuracy, with error bars included to illustrate the standard deviation of the accuracy figures.
 After meticulous analysis, the following outcomes have been derived: Training and validation accuracy undergo a gradual ascent, showcasing the model’s persistent learning and adaptation throughout the training process.
@@
 VI Summary and conclusions
-This paper focuses on applying a lightweight attention mechanism to a ResNet50 model for classifying medical histopathology tissue images, mainly when dealing with limited-scale datasets.
-The ResNet50 model incorporates the lightweight attention mechanism and a dropout layer to optimize the pre-trained deep network layer.
+This paper proposed a DALAResNet50 model for classifying medical histopathology tissue images, mainly when dealing with limited-scale imbalanced breast cancer histopathology image datasets.
+The ResNet50 model incorporates the lightweight attention mechanism to optimize the pre-trained deep network layer.
 The performance evaluation includes metrics such as accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of this paper model’s performance compared to existing techniques.
-The experimental results highlight significant advantages, particularly when addressing the challenges of unbalanced breast cancer datasets.
+The experimental results highlight significant advantages, particularly when addressing the challenges of imbalanced breast cancer datasets.
 The model demonstrates improved robustness, generalization, faster convergence during training, and effective prevention of overfitting.

### v9

@@
 The reported results of DALAResNet50 have been shown to outperform the compared approaches regarding accuracy, F1 score, IBA, and GMean, demonstrating significant robustness and broad applicability when dealing with different magnifications and imbalanced breast cancer datasets.
+Additionally, to address the interpretability challenges, we introduce the Dynamic Threshold Grad-CAM (DT Grad-CAM) method.
+DT Grad-CAM enhances the traditional Grad-CAM approach by applying adaptive thresholding using Otsu’s method, providing clearer and more focused visualizations of significant regions.
+This method improves the clarity and relevance of the highlighted areas, making it easier for medical experts to interpret the model’s predictions and understand the critical features contributing to the classification.
+The integration of DT Grad-CAM with DALAResNet50 not only ensures high classification performance but also offers enhanced interpretability, crucial for reliable and transparent breast cancer diagnosis.
 Index Terms:
@@
 Moreover, this paper presents a comparative analysis between the proposed DALAResNet50 model and the standard ResNet model, as illustrated in Table 2.
+II-D Grad-CAM-based Visualization Techniques in Histopathology Image Analysis
+Grad-CAM (Gradient-weighted Class Activation Mapping) is a widely used technique for visualizing the regions of an image that are most relevant to a neural network’s predictions.
+This technique has been extensively explored in the field of histopathology image analysis to improve model interpretability and provide insights into the decision-making process of deep learning models [59].
+In histopathology, accurate interpretation of tissue images is crucial for diagnosing diseases such as cancer.
+Traditional Grad-CAM methods generate heatmaps by computing the gradients of the target class with respect to the feature maps of the last convolutional layer, highlighting the regions of the image that contribute most to the classification decision.
+This allows pathologists to understand which areas of the tissue image the model focuses on, aiding in verifying the model’s predictions.
+However, conventional Grad-CAM has limitations, particularly in dealing with noisy activations and less distinct regions, which can hinder the clarity and precision of the visualizations.
+Various enhancements have been proposed to address these challenges.
+For instance, Guided Grad-CAM combines Grad-CAM with guided backpropagation to produce high-resolution class-discriminative visualizations, offering finer details of the important regions [60].
+Additionally, Score-CAM removes the dependency on gradients by using the target class score, improving the heatmaps’ stability and robustness[61].
+“
+To further enhance the interpretability and effectiveness of visual explanations in histopathology image analysis, we chose to implement the Dynamic Threshold Grad-CAM (DT Grad-CAM) method.
+DT Grad-CAM addresses the limitations of conventional Grad-CAM by applying adaptive thresholding using Otsu’s method, which dynamically determines the optimal threshold for highlighting significant regions.
+This approach reduces noise and enhances the clarity of the visualizations, making it easier to identify and interpret the critical features that contribute to the model’s predictions.
+By focusing on the most relevant areas and minimizing the influence of less important regions, DT Grad-CAM provides more precise and interpretable heatmaps, which are particularly beneficial in clinical settings where accurate and transparent decision-making is essential.
 III Methods
@@
 Module Name Output Size 50 Layers
-Conv1 $112\times 112112 × 112$ $7\times 77 × 7$, 64 stride=2
-[…]
-[…]
-[…]
-[…]
-Pooling $1\times 11 × 1$ Average Pool, 1000-d FC.
+Conv1 $112\times 112$ $7\times 7$, 64 stride=2
+Conv2_x $56\times 56$ $\begin{Bmatrix}1\times 1,&64\\ 3\times 3&64\\ 1\times 1&256\end{Bmatrix}$ x3
+Conv3_x $28\times 28$ $\begin{Bmatrix}1\times 1,&128\\ 3\times 3&128\\ 1\times 1&512\end{Bmatrix}$ x4
+Conv4_x $14\times 14$ $\begin{Bmatrix}1\times 1,&256\\ 3\times 3&256\\ 1\times 1&1024\end{Bmatrix}$ x6
+Conv5_x $7\times 7$ $\begin{Bmatrix}1\times 1,&512\\ 3\times 3&512\\ 1\times 1&2048\end{Bmatrix}$ x3
+Pooling $1\times 1$ Average Pool, 1000-d FC.
 Note: The term ”1000-d FC” in the pooling layer refers to a fully connected layer consisting of 1000 neurons.
@@
 III-B3 DALAResNet50 model optimization
-During the optimization phase of the DALAResNet50 model, it is adopted that focuses sequentially on pooling layers, activation layers, and nonlinear feature adjustments to enhance the model’s performance and adaptability.
+During the optimization phase of the DALAResNet50 model, a sequential focus is placed on pooling layers, activation layers, and nonlinear feature adjustments to enhance the model’s performance and adaptability.
 The related steps are as follows:
@@
 The introduction of adaptive average pooling not only simplifies the network’s complexity but also enhances the model’s ability to capture important information from different parts of the image, laying a solid foundation for subsequent feature processing.
+Specifically, the adaptive average pooling operation is defined as follows:
+    $y_{i}=\frac{1}{H_{i}\times W_{i}}\sum_{h=1}^{H_{i}}\sum_{w=1}^{W_{i}}x_{i,h,w}$ (1)
+    where $y_{i}$ is the output of the adaptive average pooling layer for the $i$-th feature map, $H_{i}$ and $W_{i}$ are the height and width of the $i$-th feature map, respectively, and $x_{i,h,w}$ represents the value at position $(h,w)$ in the $i$-th feature map.
+This approach establishes a solid foundation for subsequent feature processing, allowing the model to effectively capture and utilize global contextual information from the input images.
   2. 2.
@@
 This function is chosen for its unique capability to maintain a slight yet nonzero gradient when the unit is inactive, effectively combating the problem of vanishing gradients and significantly boosting the training efficiency of deep neural networks.
+The LeakyReLU activation function is defined as follows:
+    $f(x)=\begin{cases}x&\text{if }x>0\\ \alpha x&\text{otherwise}\end{cases}$ (2)
+    where $f(x)$ represents the LeakyReLU function and $\alpha$ is a small constant (typically 0.01) that allows a small gradient when the input $x$ is negative.
 Following this, the ReLU activation function [11] is applied in the subsequent fully connected layer to optimize model performance further and enhance the representation of image features.
+The ReLU activation function is defined as:
+    $f(x)=\max(0,x)$ (3)
+    where (f (x)) is the ReLU function, which outputs (x) if (x) is positive and 0 otherwise.
 The careful combination of LeakyReLU and ReLU activation functions dramatically improves the model’s ability to process and learn complex image characteristics, providing strong support for accurately classifying breast cancer images.
+By employing LeakyReLU in the initial phase, the model effectively addresses the vanishing gradient problem, while the subsequent use of ReLU enhances the network’s nonlinearity and expressive power, thereby improving its overall performance and robustness.
   3. 3.
@@
 This approach notably improves the model’s feature processing capability, leading to higher accuracy and efficiency in classifying breast cancer images.
+The element-wise multiplication operation is defined as follows:
+    $z_{i}=x_{i}\cdot y_{i}$ (4)
+    where $z_{i}$ is the result of the element-wise multiplication of the $i$-th feature map, and $x_{i}$ and $y_{i}$ are the corresponding elements of the feature maps being multiplied.
+This operation allows the model to adjust the features dynamically, enhancing its ability to capture complex patterns and interactions within the data.
 Fig.
@@
 These flowcharts highlight the newly added features in the DALAResNet50.
-A comprehensive comparative analysis of these differences is detailed in Table 2.
-The visualization of the DALAResNet50 approach is depicted in Fig.
-4.
+Table 2 details a comprehensive comparative analysis of these differences.
+Fig.
+4 depicts the visualization of the DALAResNet50 approach.
 IV Experiment
@@
 5 presents these pathological images as a practical example.
+Since the BreakHis dataset is a publicly available dataset, this study does not require ethical approval.
+Additionally, publicly available datasets are typically subject to stringent quality control measures; therefore, this study does not separately consider image quality issues.
 40x
@@
 RseNet50 [9] 0.927 0.860/0.927 0.914 0.914 0.944 0.891/0.944 0.922 0.922
-DALAResNet50 Model 0.985 0.970/0.985 0.980 0.980 0.987 0.974/0.987 0.985 0.985
+DALAResNet50 0.985 0.970/0.985 0.980 0.980 0.987 0.974/0.987 0.985 0.985
 Network Models 200X 400X
@@
 RseNet50 [9] 0.960 0.922/0.960 0.947 0.947 0.925 0.856/0.925 0.918 0.918
-DALAResNet50 Model 0.979 0.958/0.979 0.98 0.98 0.943 0.891/0.944 0.945 0.945
+DALAResNet50 0.979 0.958/0.979 0.98 0.98 0.943 0.891/0.944 0.945 0.945
 Note: In the table, the DenseNet-121 model indicates that employed a pre-trained DenseNet-121 model for transfer learning.
@@
 ResNet50 [9] 2920.53 2933.26 2307.94 2097.33 21797672 2564.77
-DALAResNet50 Model 2133.02 2319.50 1990.58 1896.82 25611842 2084.98
+DALAResNet50 2133.02 2319.50 1990.58 1896.82 25611842 2084.98
 Note: In the table, the DenseNet-121 model indicates that a pre-trained DenseNet-121 model was employed for transfer learning.
@@
 To emphasize the effectiveness of the proposed approach, an in-depth comparison of all these models has been conducted, achieving significant improvements in various evaluation metrics.
-40X
-100X
-200X
-400X
-Fig. 6: Training and validation loss with error bars
 IV-E Experiment metrics
@@
 This further reinforces the proposed model’s efficiency and strong potential for medical image classification, highlighting superior performance and effectiveness.
-V Performance evaluation and visualization
-To effectively monitor the model’s performance and gain insights into the training process, it is essential to concentrate on various crucial metrics, such as training loss, validation loss, training accuracy, and validation accuracy.
-Visualizing these metrics offers a comprehensive view of the model’s training behavior.
-An in-depth analysis of these key metrics and their implications is outlined below:
+IV-F Ablation study
+In this ablation study, we compared the performance of the DALAResNet50 model and the standard ResNet50 model for benign and malignant classification at 40x, 100x, 200x, and 400x magnifications.
+As shown in Table 6, the DALAResNet50 model outperformed the standard ResNet50 model across all metrics, including accuracy, precision, recall, F1 score, IBA, and GMean.
+Additionally, we provided confusion matrices for both models on the test set, illustrated in Fig.
+6.
+The results demonstrated that the DALAResNet50 model exhibited higher sensitivity and specificity at all magnifications, significantly reducing the number of false positives and false negatives.
+These findings suggest that the DALAResNet50 model surpasses the standard ResNet50 model in terms of classification accuracy and reliability.
+V DT Grad-CAM
+The interpretability of deep learning models is crucial, particularly in areas such as breast cancer diagnosis.
+Grad-CAM is a popular technique for visualizing class-specific regions of interest in images.
+This article introduces an enhancement called DT Grad-CAM (Dynamic Threshold Grad-CAM), which improves Grad-CAM by applying adaptive thresholds using the Otsu method, enhancing the visualizations’ clarity and relevance.
+V-A DT Grad-CAM method
+The DT Grad-CAM method enhances the traditional Grad-CAM approach through several steps.
+Each step improves the visualization of class-specific regions, making the results more interpretable and relevant for diagnostic purposes.
+1. Grad-CAM Calculation
+Grad-CAM uses the following formula to compute the gradient-weighted feature map:
+$L_{c}^{\text{Grad-CAM}}=\text{ReLU}\left(\sum_{k}\alpha_{k}^{c}A^{k}\right)$ (5)
+where:
+  • $L_{c}^{\text{Grad-CAM}}$ is the Grad-CAM heatmap for class $c$.
+  • $A^{k}$ is the $k$-th feature map of the convolutional layer.
+  • $\alpha_{k}^{c}$ is the weight for feature map $A^{k}$, defined as the global average pooling of the gradients of class $c$ with respect to $A^{k}$:
+$\alpha_{k}^{c}=\frac{1}{Z}\sum_{i}\sum{j}\frac{\partial y^{c}}{\partial A_{ij}^{k}}$ (6)
+where:
+  • $y^{c}$ is the score for class $c$.
+  • $\frac{\partial y^{c}}{\partial A_{ij}^{k}}$ is the gradient of the score for class $c$ with respect to the feature map $A^{k}$ at location $(i,j)$.
+  • $Z$ is the total number of elements in the feature map (i.e., width times height).
+This step leverages the gradient information to highlight the most important features contributing to the class prediction, providing an intuitive visual explanation of the model’s decision-making process.
+2. Otsu Threshold Calculation
+After generating the Grad-CAM heatmap, DT Grad-CAM applies Otsu’s method to determine the optimal threshold value. Otsu’s method calculates the threshold $T$ that maximizes the between-class variance:
+$\sigma_{B}^{2}(T)=\omega_{1}(T)\omega_{2}(T)[\mu_{1}(T)-\mu_{2}(T)]^{2}$ (7)
+where:
+  • $\sigma_{B}^{2}(T)$ is the between-class variance for threshold $T$.
+  • $\omega_{1}(T)$ and $\omega_{2}(T)$ are the probabilities of foreground and background pixels at threshold $T$, respectively.
     40x
@@
     400x
-Fig. 7: Training and validation accuracy with error bars
-V-A Loss analysis
+    Figure 6: Comparison of the confusion matrices of the DALAResNet50 model and the standard ResNet50 model.
+    40X- Original
+    40X-DTGrad-CAM
+    40X- Grad-CAM
+    100X- Original
+    100X-DTGrad-CAM
+    100X- Grad-CAM
+    200X- Original
+    200X-DTGrad-CAM
+    200X- Grad-CAM
+    400X- Original
+    400X-DTGrad-CAM
+    400X- Grad-CAM
+    Figure 7: Comparison of Original Histopathology Images with DT Grad-CAM and Grad-CAM Visualizations at Different Magnifications
+  • $\mu_{1}(T)$ and $\mu_{2}(T)$ are the means of the foreground and background pixels at threshold $T$, respectively.
+By iterating through all possible thresholds $T$, the one that maximizes $\sigma_{B}^{2}(T)$ is selected as the Otsu threshold.
+Using Otsu’s method allows for adaptive thresholding based on the image’s specific characteristics, ensuring that the most relevant regions are highlighted dynamically rather than relying on a fixed threshold.
+3. Thresholding
+The Otsu threshold is then used to binarize the Grad-CAM heatmap:
+$L_{c}^{\text{Thresh}}(x,y)=\begin{cases}L_{c}^{\text{Grad-CAM}}(x,y)&\text{if }L_{c}^{\text{Grad-CAM}}(x,y)>T\\ 0&\text{otherwise}\end{cases}$ (8)
+where:
+  • $L_{c}^{\text{Thresh}}$ is the thresholded Grad-CAM heatmap.
+  • $T$ is the optimal threshold determined by Otsu’s method.
+This thresholding step ensures that only the most significant regions in the Grad-CAM heatmap are highlighted, improving the visualization’s clarity and focus by removing less relevant areas.
+V-B Comparison between DT Grad-CAM and Grad-CAM Visualization
+Fig.
+5 illustrates a side-by-side comparison of original images, DT Grad-CAM visualizations, and traditional Grad-CAM visualizations at various magnifications (40X, 100X, 200X, and 400X).
+Each row shows the original image, the DT Grad-CAM visualization, and the Grad-CAM visualization, respectively.
+The DT Grad-CAM method consistently provides clearer and more focused representations of significant regions across all magnifications.
+By applying adaptive thresholding using Otsu’s method, DT Grad-CAM reduces noise and enhances the visibility of critical areas, which is particularly evident when comparing the middle columns (DT Grad-CAM) with the right columns (Grad-CAM).
+This adaptive approach results in sharper and more interpretable visualizations, highlighting the relevant features more distinctly and minimizing the influence of irrelevant regions.
+Consequently, DT Grad-CAM offers superior clarity and relevance, making it a valuable tool for improving model interpretability, especially in diagnostic applications such as breast cancer imaging.
+Table 8 provides a detailed comparison between DT Grad-CAM and Grad-CAM, highlighting the key differences and advantages of DT Grad-CAM.
+The adaptive thresholding in DT Grad-CAM enhances the clarity and relevance of the visualizations by minimizing noise and focusing on the most significant features.
+The clarity of visualization is significantly improved in DT Grad-CAM by isolating the key areas of interest and reducing the influence of less relevant regions.
+This enhancement in clarity directly translates to better interpretability, as DT Grad-CAM offers a sharper and more meaningful highlight of critical regions, making it easier for users to understand and interpret the model’s focus areas.
+Noise reduction is another major advantage of DT Grad-CAM, as it effectively reduces noise in the visualization, ensuring that only the most pertinent areas are highlighted.
+This is particularly useful in diagnostic applications, such as breast cancer imaging, where precise and clear visualization of important regions is crucial for accurate analysis.
+In contrast, traditional Grad-CAM lacks these adaptive enhancements, often resulting in more noise and less precise highlighting of important regions, which can obscure the interpretation and impact the accuracy of diagnostic applications.
+TABLE VIII: Comparison between DT Grad-CAM and Grad-CAM
+Feature DT Grad-CAM Grad-CAM
+Adaptive Thresholding Utilizes Otsu’s method to dynamically determine the optimal threshold, enhancing the clarity and relevance of the highlighted regions by minimizing noise and focusing on the most significant features.
+Does not employ adaptive thresholding, often leading to more noise and less precise highlighting of important regions.
+Clarity of Visualization Provides clearer and more focused visualizations by isolating the key areas of interest and reducing the influence of less relevant regions.
+Visualizations may be more dispersed and less distinct, making it harder to interpret the significant regions accurately.
+Interpretability Enhances model interpretability by offering a sharper and more meaningful highlight of critical regions, making it easier for users to understand and interpret the model’s focus areas.
+Provides a basic level of interpretability but can be less effective in highlighting the most relevant features clearly.
+Noise Reduction Effectively reduces noise in the visualization, ensuring that only the most pertinent areas are highlighted, thus improving the overall quality of the heatmap.
+Higher levels of noise can obscure the interpretation, as less important regions may also be highlighted.
+Relevance in Diagnostic Applications Particularly useful in diagnostic applications, such as breast cancer imaging, where precise and clear visualization of important regions is crucial for accurate analysis.
+Useful but may lack the precision required for detailed diagnostic purposes, potentially impacting the accuracy of interpretations.
+VI Performance evaluation and visualization
+To effectively monitor the model’s performance and gain insights into the training process, it is essential to concentrate on various crucial metrics, such as training loss, validation loss, training accuracy, and validation accuracy.
+Visualizing these metrics offers a comprehensive view of the model’s training behaviour.
+An in-depth analysis of these key metrics and their implications is outlined below:
+VI-A Loss analysis
 This study deployed a DALAResNet50 model for the classification of breast cancer.
@@
 In the later training phases, the validation loss stabilizes, suggesting that the model avoids overfitting on the validation dataset and exhibits robust generalization capabilities to unseen data.
-V-B Accuracy analysis
+VI-B Accuracy analysis
 In Fig.7, the progression of training and validation accuracy across various image resolutions (40X, 100X, 200X, and 400X) is illustrated [58].
@@
 These findings are pivotal in comprehending the model’s performance and can facilitate further optimization for enhanced breast cancer classification results.
-VI Summary and conclusions
+VII DALAResNet50’s Future Therapeutic Impact and Integration into Existing Diagnostic Processes
+DALAResNet50 model promises to significantly impact the future of breast cancer treatment by enhancing diagnostic accuracy and providing clear, interpretable visualizations through the integration of DT Grad-CAM.
+As healthcare systems increasingly adopt artificial intelligence (AI) technologies, This model can streamline diagnostic workflows, reduce diagnostic errors, and facilitate earlier detection of breast cancer, which is crucial for improving patient outcomes.
+Incorporating DALAResNet50 into existing diagnostic processes can enhance the efficiency and accuracy of histopathology evaluations.
+Pathologists can leverage the model’s high classification accuracy to identify cancerous tissues more confidently, thus supporting more accurate and timely diagnoses.
+The use of DT Grad-CAM further aids pathologists by providing visual explanations for the model’s predictions, highlighting critical regions within histopathology images, and ensuring transparency in the diagnostic process.
+Moreover, the DALAResNet50 model can serve as a valuable second-opinion tool, assisting pathologists in confirming their initial diagnoses and potentially identifying cases that may require further review.
+This integration can lead to a more robust diagnostic process, where AI augments human expertise, reducing the cognitive load on pathologists and minimizing the risk of oversight.
+As the model continues to be refined and validated with larger and more diverse datasets, its application can extend beyond breast cancer to other types of histopathological analyses.
+This scalability and adaptability make DALAResNet50 a versatile tool in the broader field of medical diagnostics.
+In the future, integrating DALAResNet50 with electronic health record (EHR) systems can further streamline patient management by providing seamless access to diagnostic results and AI-generated insights.
+This integration would facilitate multidisciplinary collaboration among healthcare providers, ensuring that treatment plans are informed by the most accurate and comprehensive diagnostic information available.
+Overall, the DALAResNet50 model, combined with DT Grad-CAM’s interpretability, holds great promise for transforming breast cancer diagnosis and treatment.
+Its integration into existing diagnostic workflows can lead to more precise, reliable, and efficient medical care, ultimately enhancing patient outcomes and advancing the field of histopathology.
+40X
+100X
+200X
+400X
+Fig. 8: Training and validation accuracy with error bars
+40X
+100X
+200X
+400X
+Fig. 9: Training and validation loss with error bars
+VIII Summary and conclusions
 This paper proposed a DALAResNet50 model for classifying medical histopathology tissue images, mainly when dealing with limited-scale imbalanced breast cancer histopathology image datasets.
 The ResNet50 model incorporates the lightweight attention mechanism to optimize the pre-trained deep network layer.
-The performance evaluation includes metrics such as accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of this paper model’s performance compared to existing techniques.
+The performance evaluation includes metrics such as accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of this paper’s model performance compared to existing techniques.
 The experimental results highlight significant advantages, particularly when addressing the challenges of imbalanced breast cancer datasets.
@@
 These findings provide a solid basis for enhancing classification accuracy.
-In future research, investigating more investigations of breast multiclassification problems are planned.
+Furthermore, this study introduces the DT Grad-CAM method to address the interpretability challenges often associated with deep learning models.
+DT Grad-CAM enhances the traditional Grad-CAM approach by applying adaptive thresholding using Otsu’s method, which provides clearer and more focused visualizations of significant regions.
+This method improves the clarity and relevance of the highlighted areas, making it easier for medical experts to interpret the model’s predictions and understand the critical features contributing to the classification.
+The integration of DT Grad-CAM with DALAResNet50 ensures high classification performance and enhanced interpretability, crucial for reliable and transparent breast cancer diagnosis.
+In future research, more investigations into breast multiclassification problems are planned.

### v10

@@
 The DALAResNet50 method was tested on breast cancer histopathology images from the BreakHis Database across magnification factors of 40X, 100X, 200X, and 400X, achieving accuracies of 98.5%, 98.7%, 97.9%, and 94.3%, respectively.
-It was also compared with established deep learning models such as SEResNet50, DenseNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50.
-The reported results of DALAResNet50 have been shown to outperform the compared approaches regarding accuracy, F1 score, IBA, and GMean, demonstrating significant robustness and broad applicability when dealing with different magnifications and imbalanced breast cancer datasets.
-Additionally, to address the interpretability challenges, we introduce the Dynamic Threshold Grad-CAM (DT Grad-CAM) method.
+It was also tested on the four-class classification BACH dataset and the three-class classification Mini-DDSM dataset, where it also performed excellently.
+Additionally, the model was compared with established deep learning models such as SEResNet50, DenseNet121, VGG16, VGG16Inception, ViT, Swin-Transformer, Dinov2_Vitb14, and ResNet50.
+The reported results of DALAResNet50 have been shown to outperform the compared approaches regarding accuracy, F1 score, IBA, and GMean, demonstrating significant robustness and broad applicability when dealing with different magnifications and imbalanced breast cancer datasets, proving the model’s wide applicability.
+To address the interpretability challenges, we introduce the Dynamic Threshold Grad-CAM (DT Grad-CAM) method.
 DT Grad-CAM enhances the traditional Grad-CAM approach by applying adaptive thresholding using Otsu’s method, providing clearer and more focused visualizations of significant regions.
@@
 =-21pt
-Figure 1: Overview of Convolutional Neural Network
 I Introduction
@@
 CNNs comprise convolutional, pooling, and fully connected layers, collaboratively facilitating the robust classification of medical images.
-The convolutional layer identifies local image features; the pooling layer reduces data dimensions while retaining critical information; and the fully connected layer maps these features to output classes, as illustrated in Fig.
-1.
+The convolutional layer identifies local image features; the pooling layer reduces data dimensions while retaining critical information; and the fully connected layer maps these features to output classes.
 Despite deep learning’s significant achievements in medical image classification, conventional CNN architectures often struggle with imbalanced medical image data.
@@
 (a) This study introduces a novel lightweight attention mechanism network that adaptively adjusts channel feature extraction to enhance feature representation.
-This approach optimizes the network’s ability to discern relevant features for breast cancer classification.
-(b) Integrating the attention mechanism with ResNet50 [9] creates a feature-rich model, achieving more efficient feature extraction by combining the strengths of both techniques.
-Specifically, within the attention mechanism’s fully connected network component, LeakyReLU [10] and ReLU [11] activation functions are employed.
-LeakyReLU is chosen for its capability to alleviate the issue of vanishing gradients; ReLU is used to improve the representation of image features, facilitating the capture of multi-scale spatial patterns more effectively.
-(c) The paper tests images at various magnifications, demonstrating the network’s scale invariance and superior performance over other contemporary methods.
-(d) The model undergoes comprehensive evaluation, focusing on analyzing loss.
-The findings indicate stable performance during the later stages of training, effectively minimizing overfitting and demonstrating strong generalization capabilities for new data.
+This method optimizes the network’s ability to recognize features relevant to breast cancer classification.
+Performance was evaluated on the BreakHis dataset for binary classification, the BACH dataset for four-class classification, and the Mini-DDSM dataset for three-class classification.
+Evaluation metrics included accuracy, F1 score, IBA, and GMean, which allow for a comprehensive assessment of the model’s performance compared to existing techniques.
+Experimental results highlighted significant advantages, especially in addressing the challenges of imbalanced breast cancer datasets.
+The model demonstrated improved robustness, generalization, faster convergence during training, and effective prevention of overfitting.
+(b) This paper introduces a DT Grad-CAM method to address the interpretability challenges commonly associated with deep learning models.
+DT Grad-CAM enhances the traditional Grad-CAM by applying adaptive thresholding using Otsu’s method, providing clearer and more focused visualizations of important regions.
+This approach improves the clarity and relative prominence of highlighted areas, making it easier for medical experts to interpret the model’s predictions and understand the key features contributing to the classification.
 Organization of the paper:
@@
 As shown in Fig.
-2, the proposed approach includes a histopathology breast cancer database and components for preprocessing and data augmentation, transfer learning based on ResNet50, and classification based on an improvised dual-activated fully connected layer.
+1, the proposed approach includes a histopathology breast cancer database and components for preprocessing and data augmentation, transfer learning based on ResNet50, and classification based on an improvised dual-activated fully connected layer.
 The proposed approach involves the following stages for breast cancer classification using histopathology images:
@@
 Note: The term ”1000-d FC” in the pooling layer refers to a fully connected layer consisting of 1000 neurons.
-TABLE II: Comparison between DALAResNet50 model and standard ResNet model
-Feature DALAResNet50 model Standard ResNet model
-Adaptive Average Pooling Adaptive average pooling addresses ”limitations in global feature extraction” by enhancing the model’s capacity to process global contextual information, significantly improving performance in tasks requiring comprehensive image understanding.
-Not available
-Channel Recalibration A two-layer fully connected network for channel recalibration optimize ”feature priority precision,” enhancing the model’s ability to identify and focus on the most relevant features for accurate classification.
-Not available
-Activation Functions The dual activation functions of LeakyReLU and ReLU address the ”inefficiency of activation functions.” By incorporating LeakyReLU, the network avoids the dying ReLU problem, as it allows a small, nonzero gradient when the unit is not active, ensuring continuous learning dynamics and preventing neurons from becoming inactive across all network layers.
-This combination maintains the model’s learning capability, potentially enhancing the training process.
-ReLU only
-Element-wise multiplication Element-wise multiplication, as denoted by ’Scale’ in Fig.
-3, enhance ”Feature Adaptability” by allowing dynamic, nonlinear adjustment of features according to a specific input, improving the network’s flexibility and utilization of information.
-Not available
 Fig. 1: The framework of DALAResNet50 model
@@
 Fig. 3: Visualization of DALAResNet50 model
+40x
+100x
+200x
+400x
+Fig. 4: Images of breast cancer pathology at four magniations
 III-B Dual-activated lightweight attention ResNet50
@@
 These flowcharts highlight the newly added features in the DALAResNet50.
-Table 2 details a comprehensive comparative analysis of these differences.
 Fig.
@@
 IV Experiment
-IV-A BreakHis Dataset
+IV-A Dataset
+IV-A1 BreakHis Dataset
 The BreakHis dataset, utilized in this study focusing on breast cancer histopathology images, was generously provided by scholar Spanha [38].
@@
 4 presents these pathological images as a practical example.
-Since the BreakHis dataset is a publicly available dataset, this study does not require ethical approval.
-Additionally, publicly available datasets are typically subject to stringent quality control measures; therefore, this study does not separately consider image quality issues.
-40x
-100x
-200x
-400x
-Fig. 5: Images of breast cancer pathology at four magniations
-TABLE III: BreakHis dataset: benign and malignant image distribution across magnification factors
+Since the BreakHis dataset is publicly available, this study does not require ethical approval.
+Additionally, publicly available datasets are typically subject to stringent quality control measures; therefore, this study does not consider image quality issues separately.
+IV-A2 BACH Dataset
+The BACH (Breast Cancer Histology Images) dataset [63] is a significant breast cancer histopathology image dataset provided by the ICIAR2018 Grand Challenge.
+This dataset includes 400 HE-stained histopathology images of breast cancer tissue, divided into four categories: Normal, Benign, In-situ carcinoma, and Invasive carcinoma, with 100 images in each category.
+All images have a resolution of 2048 × 1536 pixels, As shown in Table 3.
+The BACH dataset is widely used for training and testing machine learning and deep learning models to develop and evaluate breast cancer detection and classification algorithms.
+It provides high-quality annotation information, making it an essential resource in breast cancer research.
+IV-A3 Mini-DDSM Dataset
+The Mini-DDSM (Mini Digital Database for Screening Mammography) [62] is a reduced version of the DDSM dataset, primarily used for research and development in breast cancer screening.
+This dataset contains approximately 2000 mammography images, as shown in Table 4.
+a subset of the DDSM database that retains critical pathological features.
+The photos in the Mini-DDSM dataset are annotated with detailed information about lesions, including the location of masses and calcifications and their pathological diagnoses.
+TABLE II: BreakHis dataset: benign and malignant image distribution across magnification factors
 Class Magnification Factor
@@
 Total No of Images 1995 2081 2013 1820
-TABLE IV: Dataset split for experimental task
-Split Percentage Purpose
-Training 60% Model Training
-Validation 10% Hyperparameter Tuning
-Testing 30% Model Evaluation
-TABLE V: Experimental configuration
-Parameter Value
-Batch Size 32
-Epochs 50
-Learning Rate 0.0001
-Dropout 0.25
+TABLE III: BACH Dataset Details
+Category Number of Images Image Size
+Normal 100 2048 × 1536
+Benign 100 2048 × 1536
+In-situ carcinoma 100 2048 × 1536
+Invasive carcinoma 100 2048 × 1536
+Total 400 2048 × 1536
+TABLE IV: Detailed Information of the Mini-DDSM Dataset
+Category Number of Images Image Size / Format
+Normal Images 500 Variable
+Benign Images 700 Variable
+Malignant Images 800 Variable
+Total Images 2000 Variable
 IV-B Experimental environment
@@
 These carefully selected parameters provide a solid foundation for the experimental design and ensure the reliability and efficiency of the model throughout the study; the parameter settings are shown in Table 5.
-TABLE VI: Classification results of 9 network models
+TABLE V: Performance of Network Models on BreakHis Dataset (Two-Class Classification)
 Network Models 40X 100X
@@
 Note: In the table, the DenseNet-121 model indicates that employed a pre-trained DenseNet-121 model for transfer learning.
-TABLE VII: The count of parameters in the network model and the time it takes to converge.
+TABLE VI: Performance of Network Models on BACH Dataset (Four-Class Classification)
+Network Models Accuracy F1 IBA GMean
+SEResNet50 [39] 0.770 0.743/0.77 0.583 0.553
+DensNet121 [40] 0.703, 0.680/0.703 0.458 0.438
+VGG16 [41] 0.788 0.781 /0.788 0.652 0.637
+VGG16Inception [42] 0.727 0.864/0.727 0.650 0.654
+ViT [43] 0.746 0.734/ 0.745 0.580 0.555
+Swin-Transformer [44] 0.748 0.556/0.745 0.500 0.436
+Dinov2_Vitb14 [45] 0.856 0.730/0.855 0.500 0.353
+ResNet50 [9] 0.855 0.838 /0.855 0.664 0.635
+DALAResNet50 0.891 0.891/0.885 0.748 0.734
+TABLE VII: Performance of Network Models on Mini-DDSM Dataset (Three-Class Classification)
+Network Models Accuracy F1 IBA GMean
+SEResNet50 [39] 0.7006 0.699/0.701 0.709 0.768
+DensNet121 [40] 0.6301 , 0.627/0.63 0.642 0.716
+VGG16 [41] 0.6607 0.66 /0.661 0.672 0.736
+VGG16Inception [42] 0.7247 0.728/0.725 0.728 0.786
+ViT [43] 0.7515 0.72/ 0.752 0.580 0.555
+Swin-Transformer [44] 0.5110 0.51/ 0.511 0.52 0.619
+Dinov2_Vitb14 [45] 0.667 0.665/0.667 0.678 0.742
+ResNet50 [9] 0.6686 0.672 /0.669 0.677 0.74
+DALAResNet50 0.8070 0.806/0.807 0.818 0.851
+TABLE VIII: Parameters Count and Convergence Time of Network Models on BreakHis Dataset
 Network Models Convergence Time (seconds) Total Parameters Average Time
@@
 Note: In the table, the DenseNet-121 model indicates that a pre-trained DenseNet-121 model was employed for transfer learning.
+TABLE IX: Parameters Count and Convergence Time of Network Models on BACH Dataset
+Network Models Total Parameters Average Time
+SEResNet50 [39] 27074116 2921.67
+DensNet121 [40] 9029996 3275.27
+VGG16 [41] 134276934 3303.98
+VGG16Inception [42] 38148134 3432.98
+ViT [43] 51385352 2872.36
+Swin-Transformer [44] 86753470 3488.72
+Dinov2_Vitb14 [45] 86583556 2966.12
+ResNet50 [9] 21805868 3723.49
+DALAResNet50 24042692 2475.49
+TABLE X: Parameters Count and Convergence Time of Network Models on Mini-DDSM Dataset
+Network Models Total Parameters Average Time
+SEResNet50 [39] 27073603 7272.75
+DensNet121 [40] 9029483 6384.11
+VGG16 [41] 134272835 7589.96
+VGG16Inception [42] 38144035 8089.23
+ViT [43] 51376131 6555.35
+Swin-Transformer [44] 86746299 9715.44
+Dinov2_Vitb14 [45] 86582787 6432.41
+ResNet50 [9] 21797672 8268.68
+DALAResNet50 24040643 5990.31
 IV-D Baseline selection
@@
 This metric is in scenarios where class imbalance could significantly skew the performance metrics, ensuring that the model’s ability to classify less represented classes correctly is adequately considered in its overall [54].
-Table 6 showcases the model’s performance on the test set at various image resolutions (40X, 100X, 200X, and 400X).
-Across these resolutions, the model consistently outperforms nine other popular networks regarding accuracy, F1 score, IBA, and GMean.
-This consistency underlines the network’s significant advantage and underscores its potential for medical image classification tasks.
-Table 7 compares the model against eight others regarding parameter count and convergence speed.
-While DenseNet-121 has the fewest parameters, the proposed model achieves the shortest convergence times, exhibiting the fastest average speed across all resolutions (40X, 100X, 200X, and 400X).
-This further reinforces the proposed model’s efficiency and strong potential for medical image classification, highlighting superior performance and effectiveness.
+Tables 5, 6, and 7 present the model’s performance on the binary classification BrekHis dataset, the four-class classification BACH dataset, and the three-class classification Mini-DDSM dataset.
+In these results, the model consistently outperforms nine other popular networks in terms of accuracy, F1 score, IBA, and GMean.
+This consistency highlights the significant advantages of the network and underscores its potential in medical image classification tasks.
+The model was compared with eight other models in terms of parameter count and convergence speed on the binary classification BrekHe dataset, the four-class classification BACH dataset, and the three-class classification Mini-DDSM dataset, as shown in Tables 8, 9, and 10.
+Although DenseNet-121 has the fewest parameters, the proposed model achieved the shortest convergence time.
+This further reinforces the efficiency and robust potential of the proposed model in medical image classification, highlighting its superior performance and effectiveness.
 IV-F Ablation study
-In this ablation study, we compared the performance of the DALAResNet50 model and the standard ResNet50 model for benign and malignant classification at 40x, 100x, 200x, and 400x magnifications.
-As shown in Table 6, the DALAResNet50 model outperformed the standard ResNet50 model across all metrics, including accuracy, precision, recall, F1 score, IBA, and GMean.
-Additionally, we provided confusion matrices for both models on the test set, illustrated in Fig.
-6.
-The results demonstrated that the DALAResNet50 model exhibited higher sensitivity and specificity at all magnifications, significantly reducing the number of false positives and false negatives.
-These findings suggest that the DALAResNet50 model surpasses the standard ResNet50 model in terms of classification accuracy and reliability.
+In this ablation study, we compared the performance of the DALAResNet50 model with the standard ResNet50 model on the binary classification BrekHis dataset at 40x, 100x, 200x, and 400x resolutions, as shown in Table 5 above.
+We also provided the confusion matrices for both models on the test set, as illustrated in Fig.
+5.
+The results indicate that the DALAResNet50 model demonstrates higher sensitivity and specificity at all magnification levels, significantly reducing the number of false positives and false negatives.
+These findings suggest that the DALAResNet50 model outperforms the standard ResNet50 model in terms of classification accuracy and reliability.
 V DT Grad-CAM
@@
   • $\frac{\partial y^{c}}{\partial A_{ij}^{k}}$ is the gradient of the score for class $c$ with respect to the feature map $A^{k}$ at location $(i,j)$.
-  • $Z$ is the total number of elements in the feature map (i.e., width times height).
-This step leverages the gradient information to highlight the most important features contributing to the class prediction, providing an intuitive visual explanation of the model’s decision-making process.
-2. Otsu Threshold Calculation
-After generating the Grad-CAM heatmap, DT Grad-CAM applies Otsu’s method to determine the optimal threshold value. Otsu’s method calculates the threshold $T$ that maximizes the between-class variance:
-$\sigma_{B}^{2}(T)=\omega_{1}(T)\omega_{2}(T)[\mu_{1}(T)-\mu_{2}(T)]^{2}$ (7)
-where:
-  • $\sigma_{B}^{2}(T)$ is the between-class variance for threshold $T$.
-  • $\omega_{1}(T)$ and $\omega_{2}(T)$ are the probabilities of foreground and background pixels at threshold $T$, respectively.
     40x
@@
     Figure 6: Comparison of Original Histopathology Images with DT Grad-CAM and Grad-CAM Visualizations at Different Magnifications
+  • $Z$ is the total number of elements in the feature map (i.e., width times height).
+This step leverages the gradient information to highlight the most important features contributing to the class prediction, providing an intuitive visual explanation of the model’s decision-making process.
+2. Otsu Threshold Calculation
+After generating the Grad-CAM heatmap, DT Grad-CAM applies Otsu’s method to determine the optimal threshold value. Otsu’s method calculates the threshold $T$ that maximizes the between-class variance:
+$\sigma_{B}^{2}(T)=\omega_{1}(T)\omega_{2}(T)[\mu_{1}(T)-\mu_{2}(T)]^{2}$ (7)
+where:
+  • $\sigma_{B}^{2}(T)$ is the between-class variance for threshold $T$.
+  • $\omega_{1}(T)$ and $\omega_{2}(T)$ are the probabilities of foreground and background pixels at threshold $T$, respectively.
   • $\mu_{1}(T)$ and $\mu_{2}(T)$ are the means of the foreground and background pixels at threshold $T$, respectively.
@@
 In contrast, traditional Grad-CAM lacks these adaptive enhancements, often resulting in more noise and less precise highlighting of important regions, which can obscure the interpretation and impact the accuracy of diagnostic applications.
-TABLE VIII: Comparison between DT Grad-CAM and Grad-CAM
+TABLE XI: Comparison between DT Grad-CAM and Grad-CAM
 Feature DT Grad-CAM Grad-CAM
@@
 This integration would facilitate multidisciplinary collaboration among healthcare providers, ensuring that treatment plans are informed by the most accurate and comprehensive diagnostic information available.
-Overall, the DALAResNet50 model, combined with DT Grad-CAM’s interpretability, holds great promise for transforming breast cancer diagnosis and treatment.
-Its integration into existing diagnostic workflows can lead to more precise, reliable, and efficient medical care, ultimately enhancing patient outcomes and advancing the field of histopathology.
+Overall, the DALAResNet50 model, combined with DT Grad-CAM’s interpretability, holds great promise for transforming breast
 40X
@@
 Fig. 8: Training and validation loss with error bars
+cancer diagnosis and treatment.
+Its integration into existing diagnostic workflows can lead to more precise, reliable, and efficient medical care, ultimately enhancing patient outcomes and advancing the field of histopathology.
 VIII Summary and conclusions
@@
 The integration of DT Grad-CAM with DALAResNet50 ensures high classification performance and enhanced interpretability, crucial for reliable and transparent breast cancer diagnosis.
-In future research, more investigations into breast multiclassification problems are planned.
+In future research, we plan to apply the model presented in this paper to a broader range of medical fields.